### v1

# Efficient Generation of Smooth Paths with Curvature Guarantees by Mollification

Abstract
Most path following and trajectory tracking algorithms in mobile robotics require the desired path or trajectory to be defined by at least twice continuously differentiable functions to guarantee key properties such as global convergence, especially for nonholonomic robots like unicycles with speed constraints. Consequently, these algorithms typically exclude continuous but non-differentiable paths, such as piecewise functions. Despite this exclusion, such paths provide convenient high-level inputs for describing robot missions or behavior. While techniques such as spline interpolation or optimization-based methods are commonly used to smooth non-differentiable paths or create feasible ones from sequences of waypoints, they either can produce unnecessarily complex trajectories or are computationally expensive. In this work, we present a method to regularize non-differentiable functions and generate feasible paths through mollification. Specifically, we approximate an arbitrary path with a differentiable function that can converge to it with arbitrary precision. Additionally, we provide a systematic method for bounding the curvature of generated paths, which we demonstrate by applying it to paths resulting from linking a sequence of waypoints with segments. The proposed approach is computationally efficient, enabling real-time implementation on microcontrollers and compatibility with standard trajectory tracking and path following algorithms.
I Introduction
Path generation transforms a set of points, e.g., a finite collection of waypoints, into a smooth, continuous curve that satisfies geometric properties such as continuity, bounded curvature, and differentiability, to respect vehicle kinematic constraints, ultimately transforming a sparse set of discrete locations into a followable geometric path. Spline-based methods, including cubic B-splines [1, 2] and Bézier curves [3], ensure $C^{2}$ continuity and bounded curvature but may produce high curvature and cumbersome paths when the spacing of waypoints is irregular [4]. Gaussian process regression [5] and kernel smoothing [6] generate $C^{\infty}$-smooth approximations, trading exact waypoint passage for improved smoothness at the cost of $O(n^{3})$ computational complexity and potential numerical instability. Optimization-based approaches [7, 8] incorporate explicit curvature bounds through convex formulations, enabling real-time computation but requiring careful parameter tuning [9]. Sampling-based methods, like RRT* [10] with post-smoothing offer probabilistic completeness but may yield suboptimal curvature or other metrics. In summary, it is not surprising that the chosen method balances curvature control, computational complexity, and application requirements for interpolation versus approximation.
In this paper, we propose an inexpensive path generation method based on mollifying non-differentiable paths, e.g., piecewise functions, using mollifier functions. Mollifiers are smooth functions that, via convolution, approximate non-smooth functions arbitrarily close. They have been extensively used in partial differential equations [11] and in studying the non-vanishing of generalized Riemann zeta functions [12]. From an engineering perspective, mollifiers are used in signal and image processing to estimate the original probability distribution of variables under measurement error [13]. However, to the best of the authors’ knowledge, mollifiers have not been applied to path generation problems, despite their natural fit for robotics applications. For example, given a finite collection of ordered waypoints, it produces a path that can be arbitrarily close, in the Euclidean sense, to the segments connecting the waypoints sequentially. For this specific generated path, our method is well-suited for unicycle-like robots that travel at constant speed with heading-rate constraints; i.e., we can guarantee a maximum curvature for the generated path at the mild cost of slightly deviating from these connecting segments—an unavoidable trade-off, since the union of segments produces a non-differentiable path. Our approach is particularly appealing because it is computationally inexpensive as it can be run on small microcontrollers, and the resulting path, in its parametric representation, can be followed using path-following techniques based on modern guiding vector fields [14]. We also show how our method can generate 3D paths and guarantee several properties for generic sets of input points, including convexity preservation, enclosure of the generated path, and bounded maximum length when the input is also closed.
The article is organized as follows, Section III introduces the mollifier functions and the problem formulation. Section IV explores which geometrical and analytical aspects of the original function are affected by mollification. Convexity, concavity and quasiconvexity are treated, as well as the length of the path and the curvature. Section VI validates the theoretical findings numerically and experimentally. We end the article with some conclusions in Section VII.
II Notation
In this paper we use Lebesgue integration rather than Riemann integration to overcome some limitations with sets of measure zero and interchanging integral and limits. Therefore, we denote the Lebesgue measure in one dimension as $\lambda$, i.e., $\lambda:=\lambda_{1}$, and as $\lambda_{n}$ the $n$’th dimensional Lebesgue measure, and we consider the measure space $(X,\mathcal{B}(X),\lambda_{n})$ with $X\subseteq\mathbb{R}^{n}$, and $\mathcal{B}(X)$ is the Borel sigma-algebra of $X$; hence, when we say a function is measurable we are actually saying that it is Borel measurable. Consequently, all integrals in this paper must be thought as the Lebesgue integral with respect to the Lebesgue measure, even if the notation $\lambda$ is sometimes omitted in the integral.
For $p\in[1,\infty)$, we say that a measurable function $f$ is locally $p$-integrable, denoted $f\in L^{p}_{\mathrm{loc}}(X)$, if for each $x\in X$ there exists a neighborhood $U$ of $x$ such that $f|_{U}\in L^{p}(U)$.
We denote by $\operatorname{id}:X\to X$ the identity function defined as $\operatorname{id}(x)=x$. Then, the indicator function of a set $A$ is defined as
$\operatorname{ind}_{A}(x)=\begin{cases}1&x\in A,\\ 0&x\notin A\end{cases}.$
For any two sets $X,Y$ we denote $C(X,Y)$ as the set of continuous functions from $X$ to $Y$ and for $n\in\mathbb{N}\cup\{\infty\}$ we denote as $C^{n}(X,Y)$ the set of $n$-times continuously differentiable functions from $X$ to $Y$. For a set $A$ we denote its closure as $\overline{A}$ and the support of a real-valued function is defined as $\operatorname{supp}f:=\overline{\{x\in\operatorname{dom}f\mid f(x)\neq 0\}}$, where $\operatorname{dom}f$ is the domain of $f$.
Finally, a (parametric) path in $\mathbb{R}^{n}$ is a function $f:X\subset\mathbb{R}\to\mathbb{R}^{n}$. Writing $f=(f_{1},\dots,f_{n})$, we call $f_{i}$ the $i$’th component of the path for $i\in\{1,\dots,n\}$.
III Path generation by mollification
III-A Path generation requirements
This paper seeks an alternative approach to interpolation and optimization methods for the generation of paths from high-level inputs—such as waypoint sequences—that is computationally efficient, conceptually simple, and has a transparent physical interpretation. Furthermore, the generated paths must be feasible for mobile robots such as unicycles, avoiding unnecessarily complex trajectories. Technically, we consider the transformation of an arbitrary parametric path $f:\mathbb{R}\to\mathbb{R}^{n}$ into another parametric path that satisfies the following requirements casted as a formal problem.
Problem 1 (Path generation problem)
Let $f:\mathbb{R}\to\mathbb{R}^{n}$ be a measurable path and $\{\varepsilon_{i}\}_{i=1}^{n}$ be a collection of positive real numbers. Find a new path $T_{\varepsilon}(f):\mathbb{R}\to\mathbb{R}^{n}$ where $T_{\varepsilon}(f)=:\{T_{\varepsilon_{i}}(f)\}_{i=1}^{n}$ and $T_{\varepsilon_{i}}$ is a functional that acts on each component of $f$ such that:
  1. 1.
    $T_{\varepsilon_{i}}(f_{i})\to f_{i}$ as $\varepsilon_{i}\to 0$ in some sense of convergence.
  2. 2.
    For $p\in\mathbb{N}$ with $p\geq 2$, $T_{\varepsilon_{i}}(f_{i})\in C^{p}(\mathbb{R},\mathbb{R})$ for $i\in\{1,\dots,n\}$.
  3. 3.
    $T_{\varepsilon_{i}}(f_{i})$ is computationally simple. $\square$
The first requirement ensures that the generated path approximates $f$ arbitrarily well through a single independent tuning parameter per dimension, allowing a sequence of functions to be made arbitrarily close to the—potentially non-differentiable everywhere—original path. The second requirement guarantees feasibility for mobile robots like unicycles with speed constraints; we demonstrate how to bound the curvature when the input path consists of concatenated line segments when $n=2$. Finally, the third requirement enables real-time path generation with low-cost hardware.
III-B Mollifiers for the path generation
The solution to Problem 1 can be obtained by taking a weighted average of the points along the parametric path $f$ through convolution with a certain type of function known as a mollifier [11]. Let us recall the convolution operation.
Definition 1 (Convolution)
Let $f,g\in L^{1}(\mathbb{R}^{n})$. The convolution $f*g:\mathbb{R}^{n}\to\mathbb{R}$ is defined as
$(f*g)(x):=\int_{\mathbb{R}^{n}}f(y)g(x-y)\,\mathrm{d}\lambda_{n}(y).$
$\square$
We recall that the convolution is associative, bilinear and commutative. Let us now introduce the mollifier function.
Definition 2 (Mollifier)
Let $\varphi\in C^{\infty}(\mathbb{R}^{n},\mathbb{R})$ and for $\varepsilon>0$ define $\varphi_{\varepsilon}:=\frac{1}{\varepsilon^{n}}\varphi\circ\frac{\operatorname{id}}{\varepsilon}$. We call $\varphi$ a mollifier if it satisfies:
  1. 1.
    $\operatorname{supp}\varphi$ is compact.
  2. 2.
    $\int_{\mathbb{R}^{n}}\varphi\,\mathrm{d}\lambda_{n}=1$.
  3. 3.
    For any bounded $f\in C(\mathbb{R}^{n},\mathbb{R})$, $\lim_{\varepsilon\to 0}\int_{\mathbb{R}^{n}}f(x)\varphi_{\varepsilon}(x)\,\mathrm{d}\lambda_{n}(x)=f(0)$. $\square$
Let us present one of the most popular mollifiers since it will be used extensively in this paper.
Example 1
Let $\varphi:\mathbb{R}\to[0,\infty)$ be the function
$\varphi(x)=\begin{cases}c_{1}\exp\left(\frac{-1}{1-x^{2}}\right),&|x|<1,\\ 0,&|x|\geq 1\end{cases},$ (1)
where $c_{1}>0$ is a normalization constant that ensures $\int_{\mathbb{R}}\varphi\,\mathrm{d}\lambda=1$. Clearly $\operatorname{supp}\varphi=\overline{(-1,1)}=[-1,1]$ and $\operatorname{supp}\varphi_{\varepsilon}=[-\varepsilon,\varepsilon]$. Moreover, with a change of variables it can be seen that $\int_{\mathbb{R}}\varphi_{\varepsilon}\,\mathrm{d}\lambda=1$, and it can also be shown using the Lebesgue Dominated Convergence Theorem that as $\varepsilon\to 0$ the third property in Definition 2 holds. Figure 1 represents the function $\varphi_{\varepsilon}$ for different values of $\varepsilon>0$. $\square$
For several results in this paper, we will require the following assumption.
Assumption 1
The mollifier $\varphi$ is nonnegative and its support is the symmetric set around the origin [1,1][-1,1].
Note that this assumption is made for the sake of convenience. If the set is not symmetric around the origin, many of the presented results will be displaced but still apply. Also note that if the support of $\varphi$ is [a,a][-a,a], with $0<a\neq 1$, we can always rescale it to be [1,1][-1,1] via the parameter $\varepsilon$ as in Definition 2.
Figure 1: Representation of $\varphi_{\varepsilon}$ for different values of $\varepsilon>0$ where the mollifier $\varphi$ is defined in Example 1.
Now, let us consider a set of mollifiers $\{\varphi_{\varepsilon_{i}}:\mathbb{R}\to\mathbb{R}\}_{i=1}^{n}$, and define the generated path $F=(F_{1},\dots,F_{n})$ with $F_{i}:=f_{i}*\varphi_{\varepsilon_{i}}$ for all $i\in\{1,\dots,n\}$, and let us denote by $\varphi_{\varepsilon}^{(n)}$ the $n$-th derivative of the function $\varphi_{\varepsilon}$.
Theorem 1 ([11, Appendix C, Theorem 7])
Let $\varepsilon>0$, $f\in L_{loc}^{p}(\mathbb{R})$ with $p\in[1,\infty]$, and $\varphi$ be a mollifier; then the following three statements hold:
  1. 1.
    $F_{\varepsilon}\in C^{\infty}(\mathbb{R},\mathbb{R})$ and for any $n\in\mathbb{N}$ we have that $F_{\varepsilon}^{(n)}=(\varphi_{\varepsilon}*f)^{(n)}=(\varphi_{\varepsilon})^{(n)}*f$.
  2. 2.
    $\varphi_{\varepsilon}*f\to f$ pointwise almost everywhere as $\varepsilon\to 0$.
  3. 3.
[…]
While it may seem otherwise, Theorem 1 does not solve Problem 1 entirely. Indeed, the second requirement is satisfied straightforwardly. Regarding the computational efficiency from the third requirement, the computation of $F$ is straightforward. Indeed, because of the compact and symmetric support of the mollifier due to Assumption 1, we have that $F_{\varepsilon}(x)=\int_{[-\varepsilon,\varepsilon]}f(x-t)\varphi_{\varepsilon}(t)\,\mathrm{d}\lambda(t)$ is an inexpensive numerical operation, and note the compact integration interval. Furthermore, note that $F_{i}^{(n)}:=(f_{i}*\varphi_{\varepsilon_{i}})^{(n)}=f_{i}*\varphi_{\varepsilon_{i}}^{(n)}$, and also note that computing $F_{i}^{(n)}(x)$ does not require the existence of $f_{i}^{(n)}(x)$.
However, regarding the first requirement of Problem 1, Theorem 1 only gives us pointwise convergence and $L^{p}$ convergence. For path following or trajectory tracking algorithms we want to have a stronger notion of convergence for Problem 1, i.e., if $f\in L^{p}_{loc}(\mathbb{R})$ then for any mollifier $\varphi$, we want $\varphi_{\varepsilon}*f\to f$ as $\varepsilon\to 0$ uniformly. This is true if we require $f$ to be uniformly continuous; nonetheless, if $f$ is just continuous we have uniform convergence on compact subsets of $\mathbb{R}$. Let us finish by showing why mollifying a uniformly continuous $f$ solves the Path Generation Problem 1 with a stronger notion of convergence.
Theorem 2 (Uniform convergence)
Let $f\in L^{p}_{loc}(\mathbb{R})$ and let $\varphi:\mathbb{R}\to\mathbb{R}$ be a non negative mollifier. The following statements hold:
  1. 1.
    If $f$ is uniformly continuous then $F_{\varepsilon}\to f$ as $\varepsilon\to 0$ uniformly.
  2. 2.
    If $f$ is Lipschitz continuous then $F_{\varepsilon}\to f$ as $\varepsilon\to 0$ uniformly and $F_{\varepsilon}$ is Lipschitz continuous for any $\varepsilon>0$.
  3. 3.
    If $f$ is continuous then $F_{\varepsilon}\to f$ as $\varepsilon\to 0$ uniformly on compact subsets of $\mathbb{R}$.
Proof:
For the first statement, suppose $f$ is uniformly continuous. Fix $x\in\mathbb{R}$ and $\eta>0$. We know there exists a $\delta=\delta(\eta)>0$ such that $|f(a)-f(b)|<\eta$ whenever $|a-b|<\delta$. Choose $\varepsilon\in(0,\delta)$. Recall that $\int_{\mathbb{R}}\varphi=1$, thus $f(x)=\int_{\mathbb{R}}f(x)\varphi(t)\,\mathrm{d}\lambda(t)$. Therefore
$|F_{\varepsilon}(x)-f(x)|$ $=\left|\int_{(-\varepsilon,\varepsilon)}(f(x-t)-f(x))\varphi_{\varepsilon}(t)\,\mathrm{d}\lambda(t)\right|$
$\leq\int_{(-\varepsilon,\varepsilon)}|f(x-t)-f(x)|\varphi_{\varepsilon}(t)\,\mathrm{d}\lambda(t)<\eta,$
because $|x-t-x|=|t|\leq\varepsilon<\delta$. Since $x$ was arbitrary the statement follows.
For the second statement, since every Lipschitz continuous functions is uniformly continuous, the uniform convergence claim follows from the previous paragraph. Moreover, take $x,y\in\mathbb{R}$, $\varepsilon\in(0,\infty)$ and suppose that $f$ is Lipschitz with Lipschitz constant $K>0$.
$|F_{\varepsilon}(x)-F_{\varepsilon}(y)|$ $\leq\int_{\mathbb{R}}|f(x-t)-f(y-t)|\varphi_{\varepsilon}(t)\,\mathrm{d}\lambda(t)$
$\leq K|x-y|\int_{\mathbb{R}}\varphi_{\varepsilon}(t)\,\mathrm{d}\lambda(t)=K|x-y|,$
which proves that $F_{\varepsilon}$ is Lipschitz.
Finally, for the third statement, suppose now $f$ is continuous and take any compact set $K\subset\mathbb{R}$. Then $f$ is uniformly continuous on $K$. Thus, the same arguments as above can be followed noting that in this case $\delta$ depends on the compact set $K$. ∎
IV Key properties of the generated path
While Theorems 1 and 2 solve Problem 1, they do not provide further details about the properties of the resulting path. In this section, we characterize key properties of the generated path based on the input path. Specifically, we address questions such as: under what conditions does the output path preserve (local/quasi) convexity of the input? Does the output maintain monotonicity or other qualitative properties of the input, such as for step functions? How is the output path positioned relative to the input? How is the output path enclosed, and what is its length when the input is a closed path? We defer the analysis of the output path’s curvature to Section V, where we provide a detailed curvature analysis for the case when the input is a sequence of 2D or 3D waypoints. This curvature analysis serves as a systematic methodology that can be applied to other types of input paths.
[…]
IV-A Convexity properties
Convexity and local convexity are properties of great interest in the study of trajectory shapes. For example, if a vehicle attempts to follow a continuous but non-differentiable trajectory that resembles an inverted tent, such as the function $x\in\mathbb{R}\mapsto|x|$, will the mollified trajectory preserve this inverted tent shape? What if the property holds only locally? Intuitively, since mollification is a weighted average of the original function, the answer to the first question is affirmative. However, the answer to the second question depends on the parameter value $\varepsilon$. If the parameter is sufficiently large, the “average” of the function over the locally convex region may become negligible. We now present several propositions and counterexamples addressing these questions.
Proposition 1 (Convexity and mollification)
Let $f\in L^{1}_{loc}(\mathbb{R})$ be convex and $\varphi$ a nonnegative mollifier as in Definition 2; then $F_{\varepsilon}:=\varphi_{\varepsilon}*f$ is convex for any $\varepsilon>0$.
Proof:
Let $x,y\in\mathbb{R}$ and $\gamma\in[0,1]$. Since $\varphi_{\varepsilon}\geq 0$ we have that
$F_{\varepsilon}(\gamma x+$ $(1-\gamma)y)=\int_{\mathbb{R}}f(\gamma x+(1-\gamma)y-t)\varphi_{\varepsilon}(t)\,\mathrm{d}t$
$=\int_{\mathbb{R}}f(\gamma(x-t)+(1-\gamma)(y-t))\varphi_{\varepsilon}(t)\,\mathrm{d}t$
$\leq\int_{\mathbb{R}}\left[\gamma f(x-t)+(1-\gamma)f(y-t)\right]\varphi_{\varepsilon}(t)\,\mathrm{d}t$
$=\gamma F_{\varepsilon}(x)+(1-\gamma)F_{\varepsilon}(y).$
This property allows us to predict the shape of the mollified trajectory in advance. For example, if the trajectory to be followed resembles an inverted tent, the mollified trajectory $F_{\varepsilon}$ will also resemble an inverted tent for any $\varepsilon>0$. The question is whether local convexity is always preserved. This is false; local convexity is preserved only for sufficiently small $\varepsilon>0$, where the bound on $\varepsilon$ depends on the neighborhood in which the function is convex. We now present a proposition and a counterexample.
Proposition 2 (Local convexity and mollification)
Let $f\in L^{1}_{loc}(\mathbb{R})$ be a function that is convex in some set $(a,b)\subseteq\mathbb{R}$ with $\infty\leq a<b\leq\infty$. Let $\varphi$ satisfy Assumption 1. Then, for each $x,y\in(a,b)$ with $x<y$ there exists a $\delta=\delta(x,y)>0$ such that for all $\varepsilon\in(0,\delta)$ the function $F_{\varepsilon}:=f*\varphi_{\varepsilon}$ is convex in the set $(x,y)$.
Proof:
Let $x,y\in(a,b)$. Since $(a,b)$ is open there exists a real number $\delta>0$ such that $(x-\delta,y+\delta)\subset(a,b)$. The sets $V=(x-\delta,y+\delta)\subset(a,b)$ and $(x,y)\subset(a,b)$ are clearly open and convex. Choose $\xi,\zeta\in(x,y)$ and $\gamma\in[0,1]$. Then we have that $\gamma\xi+(1-\gamma)\zeta\in(x,y)$. Let $\varepsilon\in(0,\delta)$, and we know that
$F_{\varepsilon}(\gamma$ $\xi+(1-\gamma)\zeta)=\int_{\mathbb{R}}f(\gamma\xi+(1-\gamma)\zeta-t)\varphi_{\varepsilon}(t)\,\mathrm{d}\lambda(t).$
Note that by the selection of $\varepsilon$ and $\delta$ we have that $t\in(-\varepsilon,\varepsilon)\subset(-\delta,\delta)$. Thus, for any $t\in(-\varepsilon,\varepsilon)$, $\xi-t\in V$ and $\zeta-t\in V$. Since $V$ is convex we have that $\gamma(\xi-t)+(1-\gamma)(\zeta-t)\in V$ for all $t\in(-\varepsilon,\varepsilon)$ and $\gamma\in[0,1]$. Given that $f$ is convex in $(a,b)$ it is also convex in $V\subset(a,b)$, and noting $\varphi_{\varepsilon}\geq 0$ we can follow the steps of the proof of proposition 1 to reach
$F_{\varepsilon}(\gamma\xi+(1-\gamma)\zeta)$ $\leq\gamma F_{\varepsilon}(\xi)+(1-\gamma)F_{\varepsilon}(\zeta).$
Because $\xi,\zeta\in V$ and $\gamma\in[0,1]$ were arbitrary the proposition follows. ∎
Example 2
It is natural to ask whether Proposition 2 holds for any $\varepsilon>0$. That is, if $f$ is convex on an open set, is $F_{\varepsilon}$ also convex on that open set independently of $\varepsilon$ and the choice of non-negative mollifier? This is generally false, as demonstrated by the following counterexample. Consider the continuous function
$f(x)=\begin{cases}0,&x<0\\ x,&0\leq x\leq\frac{1}{2}\\ 1-x,&x>\frac{1}{2}\end{cases},$ (2)
and the mollifier of Example 1. The function (2) is convex on the open set (0.5,0.5)(-0.5,0.5). However, for $\varepsilon=3.2$, the mollified function is not convex and even lies below $f$ at every point in that open set, rather than above, as shown in Figure 2.
Figure 2: If $f$ is convex in an open neighbourhood there can exist a mollifier $\varphi$ as in example 1 and an $\varepsilon>0$ such that $F_{\varepsilon}:=f*\varphi$ is below $f$ in that neighbourhood. Blue line represents $f$, while the red one shows $F_{3.2}$ and the black one $F_{0.45}$.
Nevertheless, for $\varepsilon=0.45$, we have that $F_{\varepsilon}$ is convex on a neighborhood contained in (0.5,0.5)(-0.5,0.5). Thus, we can set $\delta=0.45$ in Proposition 2, and for any $\varepsilon<\delta$, $F_{\varepsilon}$ is convex in (0.5,0.5)(-0.5,0.5).
Remark 1
Note that Propositions 1 and 2 also hold when $f$ is concave instead of convex, with concavity replacing convexity throughout.
Before the following result, we need to prove that affine maps are invariant under mollification.
Proposition 3 (Affine functions and mollification)
Let $\varphi_{\varepsilon}$ with $\varepsilon>0$ be a mollifier with symmetric support around the origin, and let $a,b\in\mathbb{R}$; then $\varphi_{\varepsilon}*(a\operatorname{id}+b)=a\operatorname{id}+b$.
Proof:
Let $x\in\mathbb{R}$ and $\varepsilon>0$. Then
$(\varphi_{\varepsilon}*(a\operatorname{id}+b))(x)$ $=\int_{\mathbb{R}}(a(x-y)+b)\varphi_{\varepsilon}(y)\mathrm{d}\lambda(y)$
$=ax\int_{\mathbb{R}}\varphi_{\varepsilon}(y)\,\mathrm{d}\lambda(y)-a\int_{\mathbb{R}}y\varphi_{\varepsilon}(y)\,\mathrm{d}\lambda(y)$
$+b\int_{\mathbb{R}}\varphi_{\varepsilon}(y)\,\mathrm{d}\lambda(y)$
$=ax+b-a\int_{(-\varepsilon,\varepsilon)}y\varphi_{\varepsilon}(y)\,\mathrm{d}\lambda(y).$
The result follows noting that $y\in\mathbb{R}\mapsto y\varphi_{\varepsilon}(y)$ is an odd function that is integrated over a symmetric interval. ∎
We also need the Jensen’s inequality.
Theorem 3 (Jensen’s inequality)
[15, Theorem 1.6.2] Let $\varphi$ be a non-negative measurable function such that $\int_{\mathbb{R}}\varphi\mathrm{d}\lambda=1$, $g$ be any measurable function and $f$ be a convex function such that $\operatorname{dom}f\supset\operatorname{img}g$; then
$f\left(\int_{\mathbb{R}}g(x)\varphi(x)\mathrm{d}\lambda(x)\right)\leq\int_{\mathbb{R}}(f\circ g)(x)\varphi(x)\mathrm{d}\lambda(x).$
$\square$
Figure 3: Mollification of the function $f=|\operatorname{id}|$ with the mollifier $\varphi_{\varepsilon}$ as in Example 1 for different values of $\varepsilon$. Note that the function $f$ is convex, implying $F_{\varepsilon}$ is convex as demonstrated in proposition 1 and it is above the graph as shown in Proposition 4.
Now, we are ready to show that if $f$ is convex and $\varphi_{\varepsilon}$ is a non-negative even mollifier, then $\varphi_{\varepsilon}*f\geq f$ pointwise. For example, combined with Proposition 3, the following result implies that if our trajectory resembles an inverted tent, then the mollified function will also resemble a smoothed inverted tent but with the mollified function lying above the original path; see Figure 3 for the illustration.
Proposition 4 (If $f$ is convex then $F_{\varepsilon}$ is above $f$)
Let $\varphi_{\varepsilon}$ with $\varepsilon>0$ satisfying Assumption 1, $f$ be a convex function and define $F_{\varepsilon}:=\varphi_{\varepsilon}*f$. Then $F_{\varepsilon}\geq f$.
Proof:
Note that $\varphi_{\varepsilon}$ satisfies the conditions in Jensen’s inequality and that $f$ is convex. Let $x\in\mathbb{R}$, and note that the function $y\in\mathbb{R}\mapsto x-y$ is continuous, and therefore measurable. Applying Jensen’s inequality
$F_{\varepsilon}(x)=(\varphi_{\varepsilon}*f)(x)$ $=\int_{\mathbb{R}}f(x-y)\varphi_{\varepsilon}(y)\mathrm{d}\lambda(y)$
$\geq f\left(\int_{\mathbb{R}}(x-y)\varphi_{\varepsilon}(y)\mathrm{d}\lambda(y)\right)$
$=f((\varphi_{\varepsilon}*\operatorname{id})(x)).$
The result now follows from Proposition 3 since $x$ is arbitrary. ∎
Remark 2
Note that a similar result also holds if $f$ is concave, where clearly $F_{\varepsilon}$ would be below $f$ in that case.
The following result resembles Proposition 2, i.e., in order for the mollified function to be above the original function the parameter $\varepsilon$ has to be sufficiently small.
Proposition 5
Let $\varphi$ be even and satisfying Assumption 1, and $f:\mathbb{R}\to\mathbb{R}$ be a function that is convex in $(a,b)$. Then, for each $x,y\in(a,b)$ with $x<y$ there exists a $\delta=\delta(x,y)>0$ such that for all $\varepsilon\in(0,\delta)$ we have that $F_{\varepsilon}=\varphi_{\varepsilon}*f\geq f$ on $(x,y)$.
Proof:
Let $x,y\in(a,b)$. There exists a $\delta>0$ such that $(x-\delta,y+\delta)\subset(a,b)$. Choose $\varepsilon\in(0,\delta)$ and take $\xi\in(x,y)\subset(a,b)$. Then $(\xi-\varepsilon,\xi+\varepsilon)\subset(x-\varepsilon,y+\varepsilon)\subset(x-\delta,y+\delta)\subset(a,b)$. Since $f$ is convex in $(\xi-\varepsilon,\xi+\varepsilon)$ we can apply Jensen’s inequality leading to
$F_{\varepsilon}(\xi)$ $=(\varphi_{\varepsilon}*f)(\xi)=\int_{(-\varepsilon,\varepsilon)}f(\xi-t)\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)$
$\geq f\left(\int_{(-\varepsilon,\varepsilon)}(\xi-t)\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)\right)$
$=f((\varphi_{\varepsilon}*\operatorname{id})(\xi))=f(\xi),$
where the last equality comes from Proposition 3. Since $\xi\in(x,y)$ was arbitrary the result follows. ∎
Remark 3
Note that as it can be seen in Example 2, Proposition 5 does not hold for any $\varepsilon>0$. Indeed, also note that if the function is locally concave then we reach a similar results where the original function will be above the mollified function.
IV-B Quasiconvexity under mollification
Any convex function is quasiconvex, but the converse does not hold in general. Therefore, quasiconvexity is a weaker condition than convexity, and it is important to study how it is preserved under mollification. From an engineering standpoint, this is particularly useful for estimating the shape of the mollified curve.
Definition 3 (Quasiconvex function)
Let $S$ be a non empty convex set and $f:S\to\mathbb{R}$ be a real-valued function. We say that $f$ is quasiconvex if for all $\alpha\in\mathbb{R}$ the set
$S_{\alpha}:=\{x\in S\mid f(x)\leq\alpha\}$
is convex. An equivalent definition is that $f$ is quasiconvex if for all $x,y\in S$ and all $\gamma\in[0,1]$ we have
$f(\gamma x+(1-\gamma)y)\leq\max\{f(x),f(y)\}.\;\square$
We need the following lemma before proving that quasiconvexity is preserved under mollification.
Lemma 4
Let $f:\mathbb{R}\to\mathbb{R}$ be a quasiconvex function, $a,b\in\mathbb{R}$, and define $g:=a\operatorname{id}+b$; then, the following statements hold:
  1. 1.
    $f$ is measurable, and
  2. 2.
    $f\circ g$ is quasiconvex.
Proof:
First we prove that $f$ is measurable. Note that for any $\alpha\in\mathbb{R}$ the set $S_{\alpha}=\{x\in\mathbb{R}\mid f(x)\leq\alpha\}$ is a convex set, therefore path connected, hence connected. Since a set in $\mathbb{R}$ is connected if and only if it is an interval, then $S_{\alpha}$ is an interval. Thus, it is measurable because any interval is a Borel set.
We now prove that $f\circ g$ is quasiconvex. Let $\alpha\in\mathbb{R}$ and define the set
$S_{\alpha}:=\{x\in\mathbb{R}\mid(f\circ g)(x)\leq\alpha\}=\{x\in\mathbb{R}\mid f(ax+b)\leq\alpha\}.$
If $S_{\alpha}$ is empty then it is convex by definition, and if $S_{\alpha}$ is a singleton is also convex. Therefore, suppose that $S_{\alpha}$ consists of at least two elements. Choose $x,y\in S_{\alpha}$ and $\gamma\in[0,1]$. Then
$f(a(\gamma x+(1-\gamma)y)+b)$ $=f(\gamma(ax+b)+(1-\gamma)(ay+b))$
$\leq\max\{f(ax+b),f(ay+b)\}$
$\leq\max\{\alpha,\alpha\}$
$=\alpha.$
That is, $\gamma x+(1-\gamma)y\in S_{\alpha}$. Therefore $S_{\alpha}$ is a convex set, proving $f\circ g$ is quasiconvex. ∎
Having presented the main properties of quasiconvex functions, let us consider a representative quasiconvex case, the monotonic function. Interesting enough, the mollified function of a monotonic path, e.g., a staircase-like sequence of steps, will also be monotonic.
Proposition 6 (Monotonicity and mollification)
Let $f\in L^{1}_{loc}(\mathbb{R})$ be monotone increasing (resp. decreasing); then for any nonnegative mollifier $\varphi$ and $\varepsilon>0$ the function $F_{\varepsilon}:=(\varphi_{\varepsilon}*f)$ is monotone increasing (resp. decreasing).
Proof:
Suppose $f$ is monotone increasing. Let $x,y\in\mathbb{R}$ with $x>y$. Then for any $t\in\mathbb{R}$ we have xt>ytx-t>y-t, thus $f(x-t)\geq f(y-t)$.
$F_{\varepsilon}(x)-F_{\varepsilon}(y)$ $=\int_{\mathbb{R}}[f(x-t)-f(y-t)]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)\geq 0,$
since $\varphi_{\varepsilon}$ is positive. The proof is identical for monotone decreasing functions. ∎
Figure 4: Mollification of stair-step function created using shifted Heaviside step functions. The blue solid line represents the (discontinuous) stair-step function $f$, while the red solid line represents its mollification using the mollifier of equation (1) and $\varepsilon=0.5$. Due to proposition 6, $F_{\varepsilon}$ cannot present overshoots or oscillations.
Suppose the desired trajectory is the Heaviside step function $h:\mathbb{R}\to\mathbb{R}$ defined by $h(t)=\operatorname{ind}_{(0,\infty)}(t)$. By this result, its mollification is also monotonically increasing, so the mollified trajectory cannot exhibit overshoots or oscillations, see figure 4; this contrasts with the Gibbs phenomenon. We finally show that quasiconvexity is preserved under mollification with a nonnegative mollifier.
Theorem 5
Suppose $f:\mathbb{R}\to\mathbb{R}$ is a quasiconvex function and let $\varphi$ be a mollifier satisfying 1; then, for all $\varepsilon>0$ the function $F_{\varepsilon}:=f*\varphi_{\varepsilon}$ is quasiconvex.
Proof:
We are going to proceed by contradiction. Let $f$ be quasiconvex, and suppose that there exists an $\varepsilon>0$ such that $F_{\varepsilon}$ is not quasiconvex. This implies there exist $x,y\in\mathbb{R}$ and $\gamma\in[0,1]$ such that
$F(\gamma x+(1-\gamma)y)>\max\{F(x),F(y)\}.$
In particular
$F_{\varepsilon}(\gamma x+(1-\gamma)y)-F_{\varepsilon}(x)$
$=\int_{(-\varepsilon,\varepsilon)}\left[f(\gamma x+(1-\gamma)y-t)-f(x-t)\right]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)>0.$
First, we claim that the set $V_{1}:=\{t\in(-\varepsilon,\varepsilon)\mid f(\gamma x+(1-\gamma)y-t)-f(x-t)>0\}$ is measurable and it has positive measure. By Lemma 4 we know that the composition of a quasiconvex function with an affine mapping is quasiconvex, hence measurable. Since the linear combination of two measurable function results in a measurable function, the set $V_{1}$ is measurable. Thus, if $\lambda(V_{1})$ were zero, then it would negligible in integration, which would imply that
$0$ $<\int_{(-\varepsilon,\varepsilon)}\left[f(\gamma x+(1-\gamma)y-t)-f(x-t)\right]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)$
$=\int_{\left[(-\varepsilon,\varepsilon)\setminus V_{1}\right]\cup V_{1}}\left[f(\gamma x+(1-\gamma)y-t)-f(x-t)\right]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)$
$=\int_{(-\varepsilon,\varepsilon)\setminus V_{1}}\left[f(\gamma x+(1-\gamma)y-t)-f(x-t)\right]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t),$
but $(-\varepsilon,\varepsilon)\setminus V_{1}=\{t\in(-\varepsilon,\varepsilon)\mid t\notin V_{1}\}=\{t\in(-\varepsilon,\varepsilon)\mid f(\gamma x+(1-\gamma)y-t)-f(x-t)\leq 0\}$ and $\varphi$ is nonnegative, thus
$\int_{(-\varepsilon,\varepsilon)\setminus V_{1}}\left[f(\gamma x+(1-\gamma)y-t)-f(x-t)\right]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)\leq 0.$
which leads to the first contradiction. Therefore $\lambda(V_{1})>0\Longrightarrow V_{1}\neq\emptyset$. Moreover, since $f$ is quasiconvex we have that for any $t\in V_{1}$
$f(x-t)$ $<f(\gamma x+(1-\gamma)y-t)$
$=f(\gamma(x-t)+(1-\gamma)(y-t))$
$\leq\max\{f(x-t),f(y-t)\}.$
Now let us consider that
$f(y-t)>f(x-t),\quad\forall t\in V_{1}.$ (3)
because if not, we would reach a contradiction; therefore, proving that our assumption about $F_{\varepsilon}$ is false, thus proving the theorem.
The same procedure can be done considering the other point, $y$, and $F_{\varepsilon}(\gamma x+(1-\gamma)y)-F_{\varepsilon}(y)>0$, i.e., there exists a set of positive measure $V_{2}\subset(-\varepsilon,\varepsilon)$ such that
$f(\gamma x+(1-\gamma)y-t)-f(y-t)>0,\quad\forall t\in V_{2},$
and since $f$ is quasiconvex, with the same arguments as above
$f(x-t)>f(y-t),\quad\forall t\in V_{2}.$ (4)
Suppose $V_{1}\cap V_{2}\neq\emptyset$. This implies that there exists a $t\in V_{1}\cap V_{2}$ such that f(yt)<f(xt)<f(yt)f(y-t)<f(x-t)<f(y-t) which is again a contradiction, thus proving the theorem. So we just need to prove that we reach a contradiction in the case $V_{1}\cap V_{2}=\emptyset$. Take $t_{1}\in V_{1}$ and $t_{2}\in V_{2}$ and any $\beta\in[0,1]$, then
$f(y-\beta t_{1}-(1-\beta)t_{2})$ $=f(\beta(y-t_{1})+(1-\beta)(y-t_{2}))$
$\leq\max\{f(y-t_{1}),f(y-t_{2})\}$
However, $f(y-t_{2})<f(x-t_{2})$ by equation (4), thus
$f(y-\beta t_{1}-(1-\beta)t_{2})<\max\{f(y-t_{1}),f(x-t_{2})\},$
$\quad\forall\beta\in[0,1].$
Using the same approach and considering equation (3) we have that
$f(x-\beta t_{1}-(1-\beta)t_{2})<\max\{f(y-t_{1}),f(x-t_{2})\},$
$\quad\forall\beta\in[0,1].$
Since this is independent of the value of $\beta$, and in particular for $\beta=1$ and $\beta=0$ we have that
$f(y-t_{1})<\max\{f(y-t_{1}),f(x-t_{2})\}$
$f(x-t_{2})<\max\{f(y-t_{1}),f(x-t_{2})\},$
but this leads to a contradiction. Therefore, the assumption that there exists an $\varepsilon>0$ for which $F_{\varepsilon}$ is not quasiconvex is false. That is, $F_{\varepsilon}$ is quasiconvex for any $\varepsilon>0$. ∎
Remark 4
The converse is not in general true. That is, having $F_{\varepsilon}$ quasiconvex for some $\varepsilon>0$ does not imply that $f$ is quasiconvex. For example, consider the following measurable function that is not quasiconvex
$f(x)=\begin{cases}1,&x\in\mathbb{R}\setminus\mathbb{Q}\\ 2,&x\in\mathbb{Q}\end{cases},$
because the set $S_{1.5}=\{x\in\mathbb{R}\mid f(x)\leq 1.5\}=\mathbb{R}\setminus\mathbb{Q}$ is disconnected, hence it is not convex. However, since $\lambda(\mathbb{Q})=0$, we have that for any $x\in\mathbb{R}$
$F_{\varepsilon}(x)$ $=\int_{\mathbb{R}}\varphi(x-t)f(t)\mathrm{d}\lambda(t)=\int_{\mathbb{R}\setminus\mathbb{Q}}\varphi(x-t)\mathrm{d}\lambda(t)$
$=\int_{(\mathbb{R}\setminus\mathbb{Q})\cup\mathbb{Q}}\varphi(x-t)\mathrm{d}\lambda(t)=\int_{\mathbb{R}}\varphi\mathrm{d}\lambda=1,$
which is convex; thus quasiconvex too. Moreover, it is clear that if $F_{\varepsilon}$ is quasiconvex for all $\varepsilon>0$ then $f$ is quasiconvex since the pointwise limit of a family of quasiconvex functions can be shown to be quasiconvex function as well.
IV-C Closed paths
So far, we have worked with real-valued functions because we represent desired paths or trajectories as parametric functions, i.e., functions $f:\mathbb{R}\to\mathbb{R}^{n}$. Nevertheless, as previously noted, our results can be extended to functions $\phi:\mathbb{R}^{n}\to\mathbb{R}$, for which convexity-like properties can also be defined. This extension is of practical interest, since in some situations the desired path is a set of the form
$\mathcal{P}=\{x\in\mathbb{R}^{n}\mid\phi_{i}(x)=0,\quad i\in\{1,\dots,n-1\}\},$
for which previous results applying to each $\phi_{i}$. Nevertheless, there are situations in which the path arises from a parametric curve with compact domain, say $f:[a,b]\to\mathbb{R}^{n}$ where $\infty<a<b<\infty$, and it is preferable to work with $f$ as a whole rather than component by component. Since we lose some order-like properties in $\mathbb{R}^{n}$ relative to $\mathbb{R}$, we cannot directly extend convexity-like properties [^1] We refer to pure convexity properties.. Nonetheless, we still want to characterize (in advance) how mollification affects the complete function $f$. We now address the following question: given the original trajectory, does there exist a subset $U$ of $\mathbb{R}^{n}$ such that the mollified trajectory is contained in $U$ for any value of its parameter? The answer is affirmative, with $U$ being the convex hull of $f([a,b])$.
Definition 4
Let $A\subset\mathbb{R}^{n}$ be a set. Its convex hull, denoted as $\operatorname{co}(A)$ is defined as the smallest convex set that contains $A$, that is, $A\subset\operatorname{co}(A)$. $\square$
We first present a result when $\operatorname{dom}f=\mathbb{R}$.
Theorem 6
Let $f:\mathbb{R}\to\mathbb{R}^{n}$ be a measurable function and $\varphi$ be a nonnegative mollifier. Define for $t\in\mathbb{R}$ and $\varepsilon>0$
$F_{\varepsilon}(t)=(f*\varphi_{\varepsilon})(t)=\left((f_{1}*\varphi_{\varepsilon})(t),\dots,(f_{n}*\varphi_{\varepsilon})(t)\right).$
Then, given $\varepsilon>0$, we have that
$\left\{F_{\varepsilon}(t)\mid t\in\operatorname{dom}f\right\}\subset\operatorname{co}\left\{f(t)\mid t\in\operatorname{dom}f\right\}.$
Proof:
Let $U=\operatorname{co}\{f(t)\mid t\in\operatorname{dom}f\}$. Define the extended real valued function $I_{U}:\mathbb{R}^{n}\to\mathbb{R}\cup\{-\infty,\infty\}$ as
$I_{U}(x)=\begin{cases}+\infty,&x\notin U\\ 0,&x\in U\end{cases}.$
The function $I_{U}$ is clearly convex. Fix $t\in\operatorname{dom}f$. Noting that $\int_{\mathbb{R}}\varphi_{\varepsilon}\mathrm{d}\lambda=1$ and $\varphi_{\varepsilon}\geq 0$, we can apply Jensens’ Inequality in higher dimensions to get
$0\leq I_{U}(F_{\varepsilon}(t))$ $=I_{U}\left(\int_{\operatorname{supp}\varphi_{\varepsilon}}f(t-s)\varphi_{\varepsilon}(s)\mathrm{d}\lambda(s)\right)$
$\leq\int_{\operatorname{supp}\varphi_{\varepsilon}}I_{U}(f(t-s))\varphi_{\varepsilon}(s)\mathrm{d}\lambda(s).$
However, note that $I_{U}(f(t-s))=0$ for any $t-s\in\operatorname{dom}f$, and since $\operatorname{dom}f=\mathbb{R}$ then
$0\leq I_{U}(F_{\varepsilon}(t))\leq\int_{\operatorname{supp}\varphi_{\varepsilon}}I_{U}(f(t-s))\varphi_{\varepsilon}(s)\mathrm{d}\lambda(s)=0,$
i.e., $I_{U}(F_{\varepsilon}(t))=0$ so $F_{\varepsilon}(t)\in U$. Since $t$ and $\varepsilon$ are arbitrary, the claim follows. ∎
Remark 5
As previously mentioned, it is common that the path is a continuous function defined in a compact subset of $\mathbb{R}$, i.e., $f:[a,b]\to\mathbb{R}^{n}$ with $\infty<a<b<\infty$. In such a case, we can extend the function $f$ to $\mathbb{R}$ as follows to get a new continuous function
$\bar{f}(t)=\begin{cases}f(a),&-\infty<t\leq a\\ f(x),&a\leq t\leq b\\ f(b),&b\leq t<\infty\end{cases}.$
Note that $\bar{f}([a,b])=f([a,b])$ and $\bar{f}((-\infty,a]\cup[b,\infty))=\{f(a),f(b)\}$, so $\bar{f}(\mathbb{R})=f([a,b])$ and then $\operatorname{co}\bar{f}(\mathbb{R})=\operatorname{co}f([a,b])$. Then we can use as our path $\bar{f}$ instead of $f$, obtaining the result of the previous theorem, and later restricting the domain of the mollified function to $[a,b]$ again, i.e., we let $\bar{F}_{\varepsilon}=\bar{f}*\varphi_{\varepsilon}$ and use the mollified curve $F_{\varepsilon}=\bar{F}_{\varepsilon}|_{[a,b]}$, thus
$F_{\varepsilon}([a,b])\subset\bar{F}(\mathbb{R})\subset\operatorname{co}\bar{f}(\mathbb{R})=\operatorname{co}f([a,b]).$
Clearly for $t\in[a+\varepsilon,b-\varepsilon]$, $F_{\varepsilon}$ coincides with the mollification of $f$, and in $a,a+\varepsilon)$ and $(b-\varepsilon,b]$ it belongs to the convex hull of $f([a,b])$.
Having characterized the space in which the mollified path is enclosed, we now consider the relationship between the length of the original path and its mollification. First, we introduce the definition of path length for paths that do not need to be differentiable.
Definition 5 (Length of $f$)
Let $f:[a,b]\to\mathbb{R}^{n}$ be a continuous function, and $||\cdot||:\mathbb{R}^{n}\to[0,\infty)$ be any norm in $\mathbb{R}^{n}$. Let a finite set $P=\{x_{0},x_{1},\dots,x_{N}\}$, where $a=x_{0}<x_{1}<\dots<x_{N}=b$ be a partition of $[a,b]$. Then, the length of $f$ is
[…]
Note that when working with trajectories with compact domain, we must extend them as done in Remark 5.
Lemma 7
Let $f:[a,b]\to\mathbb{R}^{n}$ be a continuous function and fix $\varepsilon>0$. Let $\bar{f}:[a-\varepsilon,b+\varepsilon]$ be its continuous extension as done in remark 5. The following two statements are true:
  1. 1.
    $L(\bar{f})=L(f)$.
  2. 2.
    If $|t|\leq\varepsilon$ and $g(s)=\bar{f}(s-t)$ for all $s\in[a,b]$, then $L(g)\leq L(\bar{f})=L(f)$.
Proof:
We prove each statement separately.
  1. 1.
    Take a partition $P$ of $a-\varepsilon,b+\varepsilon]$ with $N$ elements, such that there exists $0<J<K<N$ such that $x_{j}=a$ and $x_{k}=b$. Then
[…]
    where the last inequality comes from the fact that $\{x_{J},\dots,x_{K}\}$ is a partition of $[a,b]$. Therefore, by definition of the supremum $L(\bar{f})\leq L(f)$. The inequality $L(f)\leq L(\bar{f})$ holds trivially by noting that a partition of $[a,b]$ can be extended to create a partition of $a-\varepsilon,a+\varepsilon]$ and we are summing positive terms. Therefore $L(f)=L(\bar{f})$.
  2. 2.
    Let $|t|\leq\varepsilon$ and consider a partition $P=\{x_{0},\dots,x_{N}\}$ of $[a,b]$. Clearly $P-\{t\}=\{x_{0}-t,\dots,x_{n}-t\}$ could be considered as a subset of a partition of $a-\varepsilon,a+\varepsilon]$. Therefore by constructing $P^{\prime}=(P-\{t\})\cup\{a-\varepsilon,b+\varepsilon\}$ then
[…]
    Since the supremum is the least upper bound, it follows that $L(g)\leq L(\bar{f})=L(f)$.
Now we are ready for the main result regarding the length of the generated mollified path being shorter or equal than the original.
Theorem 8
Let $f:[a,b]\to\mathbb{R}^{n}$ be a continuous function and let $\varphi$ be a nonnegative mollifier. Fixed $\varepsilon>0$ let $\bar{f}:[a-\varepsilon,b+\varepsilon]\to\mathbb{R}^{n}$ be the continuous extension as in remark 5. Define $F:[a,b]\to\mathbb{R}^{n}$ as $F=\bar{f}*\varphi_{\varepsilon}$; then $L(F)\leq L(f)$.
Proof:
Take a partition $P$ of $[a,b]$, then
[…]
thus, by definition of the supremum, $L(F)\leq L(f)$. ∎
Remark 6
Note that this does not imply that if $f$ is a geodesic between $f(a)$ and $f(b)$, then $F$ is also a geodesic between these points. This is because $F(a)\neq f(a)$ or $F(b)\neq f(b)$ may occur. While Theorem 6 and Remark 5 guarantee that $F([a,b])=(\bar{f}*\varphi_{\varepsilon})([a,b])\subset\operatorname{co}f([a,b])$, we cannot ensure that $F$ has the same starting and ending points as $f$. What Theorem 8 establishes is that by considering the actual starting and ending points of $F$, we can ensure that $L(F)\leq L(f)$. An example of this property is shown in figure 5. The original function, which is a linear interpolation of three points in $\mathbb{R}^{2}$ and whose domain is $[a,b]=[0,2]$, is extended to the domain $\varepsilon,2+\varepsilon]$ with $\varepsilon=0.5$ using the extension presented in remark 5. As it can be seen from each of its components, $F_{\varepsilon}(a)\neq f(a)$ and $F_{\varepsilon}(b)\neq f(b)$, and clearly $L(F_{\varepsilon})\leq L(f)$ as theorem 8 states. Finally note that, while $f_{2}$ can be considered a geodesic between the points $(0,0)$ and $(2,2)$ in $\mathbb{R}^{2}$ using the Euclidean norm, $(f_{2}*\varphi_{\varepsilon})$ is not a geodesic between those two points.
Figure 5: Visual representation of remark 6. The left picture represents the original path $f$ as a blue solid line, and as a red solid line its mollification $F_{\varepsilon}=(f_{1}*\varphi_{\varepsilon},f_{2}*\varphi_{\varepsilon})$ where $\varphi$ is defined in equation (1) and $\varepsilon=0.5$. The middle picture represents the same information but for the first component of the function and its mollification, while the right picture represents the same information but for the second component.
IV-D The effect of reparametrization and mollification
Suppose the desired path is encoded using a continuous function $f:[a,b]\to\mathbb{R}^{n}$ and for a given $\varepsilon>0$ we consider its mollification with parameter $\varepsilon$. As we have seen, we first need to extend the function $f$ to $\bar{f}:[a-\varepsilon,b+\varepsilon]$ as is done in Remark 5 or in Theorem 8. It may be of interest to reparametrize the curve so it is normalized, i.e., to find a function $g:[-\varepsilon,1+\varepsilon]\to[a-\varepsilon,b+\varepsilon]$, and consider $\bar{f}\circ g*\varphi_{\varepsilon}:[0,1]\to\mathbb{R}^{n}$, as the desired mollified path. How does the parameter of the mollification change under these conditions? That is, does there exist a $\eta=\eta(\varepsilon)>0$ such that
$(\bar{f}\circ g*\varphi_{\varepsilon})([0,1])=(\bar{f}*\varphi_{\eta})([a,b])?$
What we require is a continuous function $g:[-\varepsilon,1+\varepsilon]\to[a-\varepsilon,b+\varepsilon]$ that is strictly increasing and satisfies $g([0,1])=[a,b]$. However, it is easy to see that such a function must be nonlinear. In the best-case scenario where such a function exists it is invertible, if it happens to be differentiable, we arrive at the following conclusions. Suppose $g:[-\varepsilon,1+\varepsilon]\to[a-\varepsilon,b+\varepsilon]$ is a continuously differentiable, increasing function that is nonlinear but satisfies g([0,1])=[a,b]g([0,1])=[a,b]—as previously required. In this case, let $s\in[0,1]$ and applying the change of variables v(t)=g(st)v(t)=g(s-t), we get
$(f*\varphi_{\varepsilon})(s)=\int_{[-\varepsilon,\varepsilon]}f(g(s-t))\varphi_{\varepsilon}(t)\mathrm{d}t$
$=\int_{[g(s-\varepsilon),g(s+\varepsilon)]}f(v)\varphi\left(\frac{s-g^{-1}(v)}{\varepsilon}\right)\frac{1}{\varepsilon g^{\prime}(g^{-1}(v))}\mathrm{d}v$
$=\int_{[g(s-\varepsilon),g(s+\varepsilon)]}f(v)\varphi_{\varepsilon}(s-g^{-1}(v))\frac{1}{g^{\prime}(g^{-1}(v))}\mathrm{d}v.$
[…]
V Curvature guarantees from a sequence of waypoints
In this section, we show how to systematically analyze the curvature of the generated path. In particular, we provide a formula to upper bound the curvature of the mollification of a sequence of 2D or 3D waypoints connected by straight line segments, i.e., via linear interpolation. First, we restrict ourselves to the simpler case of two segments. From now on, $\mathbb{R}^{n}$ denotes either $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$.
V-A The case of three points forming two segments
Suppose the desired path can be encoded using a parametric function of the following form.
Definition 6 (Two line segments function)
Let $P_{0},P_{1},P_{2}\in\mathbb{R}^{n}$, and let $f:[0,2]\to\mathbb{R}^{n}$ be
$f(t)=\begin{cases}P_{0}+(P_{1}-P_{0})t,&t\in[0,1],\\ P_{1}+(P_{2}-P_{1})(t-1),&t\in[1,2]\end{cases}.$ (5)
We call $f$ the two-line segments function. And we call
$\bar{f}(t)=\begin{cases}P_{0}+(P_{1}-P_{0})t,&t\in(-\infty,1],\\ P_{1}+(P_{2}-P_{1})(t-1),&t\in[1,\infty)\end{cases},$ (6)
the two-lines segment extended function. Note that $\bar{f}\mid_{[0,2]}=f$. $\square$
It is straightforward to note that $f$ is differentiable in $(0,2)\setminus\{1\}$ and $\bar{f}$ in $\mathbb{R}\setminus\{1\}$. Note that in both cases, the set of points on which the functions are not differentiable form a set of measure zero, and the expressions of their derivatives are constant functions. Let $t\in\mathbb{R}\setminus\{1\}$, then
$\bar{f}^{\prime}(t)=\begin{cases}P_{1}-P_{0},&t\in(-\infty,1)\\ P_{2}-P_{1},&t\in(1,\infty)\end{cases}.$
For $r\in\mathbb{N}$ we define $\tilde{P}_{r}:=P_{r}-P_{r-1}$.
Note that both $f$ and $\bar{f}$ are continuous functions, hence locally integrable, and from Theorem 1 we have that if $F_{\varepsilon}:=\bar{f}*\varphi_{\varepsilon}$ where $\varphi$ is any mollifier and $\varepsilon>0$ then $F_{\varepsilon}^{\prime}=\bar{f}*\varphi_{\varepsilon}^{\prime}$. Fix $t\in\mathbb{R}$, and note that
$\int_{\mathbb{R}\setminus\{t-1\}}\bar{f}^{\prime}(t-s)\varphi_{\varepsilon}(s)\,\mathrm{d}s$ $=\int_{\mathbb{R}\setminus\{1\}}\bar{f}^{\prime}(s)\varphi_{\varepsilon}(s)\,\mathrm{d}s$
$=\int_{\mathbb{R}}\bar{f}^{\prime}(s)\varphi_{\varepsilon}(s)\,\mathrm{d}s,$
because $\lambda(\{1\})=0$, so we could take any function $g$ that is equal to $\bar{f}^{\prime}$ in $\mathbb{R}\setminus\{1\}$ but is defined in $t=1$, and we would reach that $\int_{\mathbb{R}}g(t-s)\varphi_{\varepsilon}(s)=\int_{\mathbb{R}\setminus\{t-1\}}\bar{f}^{\prime}(t-s)\varphi_{\varepsilon}(s)\,\mathrm{d}s$. That is, we can evaluate $(\bar{f}^{\prime}*\varphi_{\varepsilon})(1)$ even if $\bar{f}^{\prime}(1)$ is not defined. Exploiting the properties of the compact support of the mollifiers
$\int_{\mathbb{R}}\bar{f}^{\prime}(s)\varphi_{\varepsilon}(t-s)\,\mathrm{d}s$
$=\bar{f}(s)\varphi_{\varepsilon}(t-s)\mid_{s=t-\varepsilon}^{s=t+\varepsilon}-\int_{\mathbb{R}}\bar{f}(s)(-\varphi_{\varepsilon}^{\prime}(t-s))\,\mathrm{d}s$
$=\int_{\mathbb{R}}\bar{f}(s)\varphi_{\varepsilon}^{\prime}(t-s)\,\mathrm{d}s=(\bar{f}*\varphi_{\varepsilon}^{\prime})(t)=F_{\varepsilon}^{\prime}(t).$
This proves that in this case $(\bar{f}^{\prime}*\varphi_{\varepsilon})=(\bar{f}*\varphi_{\varepsilon}^{\prime})=F_{\varepsilon}^{\prime}$, which is a powerful result given that the derivative of $\bar{f}$ is constant almost everywhere.
Theorem 9
Let $f:[a,b]\to\mathbb{R}^{n}$ and consider any continuous extension of $f$ to $\mathbb{R}$, i.e., a function $\bar{f}:\mathbb{R}\to\mathbb{R}^{n}$ that is continuous, $\bar{f}\mid_{[a,b]}=f$ and it is differentiable almost everywhere in $\mathbb{R}$. If it happens that $f$ is differentiable almost everywhere in $[a,b]$ then it holds that for any $\varepsilon>0$ and mollifier
[…]
Proof:
The proof is immediate following the steps from the previous discussion. For a more general result involving distributions see [16, Sections 6.13 and 6.14]. ∎
Corollary 1
Let $f$ be the two lines segment function of equation (5) and $\bar{f}$ its extended version defined in equation (6). Then, for any $\varepsilon>0$ and any mollifier $\varphi$ it holds that
[…]
We can now exploit the results of Corollary 1 to obtain a formula for the curvature, as well as an upper bound.
V-A1 Computing the exact curvature
We know by the Corollary 1 that $F_{\varepsilon}^{\prime}=(\bar{f}*\varphi^{\prime}_{\varepsilon})=(\bar{f}^{\prime}*\varphi_{\varepsilon})$ everywhere. Consider a mollifier $\varphi$ and let $\varepsilon>0$. Given $t\in\mathbb{R}\setminus\{1\}$, note that
$\bar{f}^{\prime}(t)=\tilde{P}_{1}\operatorname{ind}_{(-\infty,1)}(t)+\tilde{P}_{2}\operatorname{ind}_{(1,\infty)}(t),$
hence for $t\in\mathbb{R}$
$F_{\varepsilon}^{\prime}(t)$ $=\int_{\mathbb{R}}\varphi_{\varepsilon}(t-s)\bar{f}^{\prime}(s)\mathrm{d}s$
$=\int_{(-\infty,1]}\varphi_{\varepsilon}(t-s)\tilde{P}_{1}\mathrm{d}s+\int_{[1,\infty)}\varphi_{\varepsilon}(t-s)\tilde{P}_{2}\mathrm{d}s$
$=\tilde{P}_{1}\int_{(-\infty,1]}\varphi_{\varepsilon}(t-s)\mathrm{d}s+\tilde{P}_{2}\int_{[1,\infty)}\varphi_{\varepsilon}(t-s)\mathrm{d}s.$
Note that if $\Phi_{\varepsilon}:\mathbb{R}\to\mathbb{R}$ is such that $\Phi_{\varepsilon}^{\prime}=\varphi_{\varepsilon}$, then
$\frac{d}{dt}\int_{[a,b]}\varphi_{\varepsilon}(t-s)\mathrm{d}s$ $=\frac{d}{dt}\int_{[t-b,t-a]}\varphi_{\varepsilon}(u)\mathrm{d}u$
$=\frac{d}{dt}(\Phi_{\varepsilon}(t-b)-\Phi_{\varepsilon}(t-a))$
$=\varphi_{\varepsilon}(t-b)-\varphi_{\varepsilon}(t-a),$
thus
$F_{\varepsilon}^{\prime\prime}(t)=\varphi_{\varepsilon}(t-1)(\tilde{P}_{1}-\tilde{P}_{2}).$
Now define
$A_{1}(t)$ $:=\int_{(-\infty,1]}\varphi_{\varepsilon}(t-s)\mathrm{d}s$
$A_{2}(t)$ $:=\int_{[1,\infty)}\varphi_{\varepsilon}(t-s)\mathrm{d}s,$
therefore, if $\kappa:\mathbb{R}\to\mathbb{R}$ is the curvature,
[…]
and noting that, due to the properties of the mollifier, $A_{1}(t)+A_{2}(t)=1$ and $A_{1}(t),A_{2}(t)\geq 0$ for all $t\in\mathbb{R}$, we have that
[…]
Equation (7) is an exact formula for the curvature at each $t\in\mathbb{R}$.
V-A2 Upper bounding the curvature
[…]
$||F_{\varepsilon}^{\prime}(t)||_{2}^{2}\geq\min_{s\in[0,1]}||s\tilde{P}_{1}+(1-s)\tilde{P}_{2}||_{2}^{2}=:\min_{s\in[0,1]}g(s).$
Note that $g$ is a differentiable convex function, so its minimum exists in the compact set $[0,1]$ and by the KKT conditions it is necessary and sufficient to find an $\bar{s}\in[0,1]$ such that $g^{\prime}(\bar{s})=0$. In this case
[…]
From which it follows that if
[…]
That is, we have found an upper bound on the curvature for two segments that is independent of $t$. Figures 6 and 7 illustrate this upper bound for two different curves. Note that when the segments have similar lengths, the upper bound equals the maximum curvature—making it the tightest possible bound. However, when one segment is significantly longer than the other, the zone of maximum curvature shifts from $t=1$ due to the mollification process. In any case, we can confidently assert that this upper bound is a good approximation of the maximum curvature, and in many cases optimal.
Figure 6: Left plot represents of a three-point-two-segment function and its mollification. Right plot represents the curvature of the function and its upper bound $U_{\kappa}$ which is the right hand side of equation (8). In this case the zone of maximum curvature corresponds to a point really close (or equal) to $t=1$. We have used the mollifier presented in Example 1.
Figure 7: Left plot represents of a three-point-two-segment function and its mollification. Right plot represents the curvature of the function and its upper bound $U_{\kappa}$ which is the right hand side of equation (8). In this case the zone of maximum curvature does not correspond to the join of the two segments. We have used the mollifier presented in Example 1.
V-B The general case
We are going to present how, the natural generalization of the previous computations to a $p>2$ segments curve, gives, in general, a worse result than considering the curvature of each pair of segments locally, and then choosing the most restrictive $\varepsilon_{max}>0$ so that all curvatures are constraint under this $\varepsilon_{max}>0$. Nevertheless, due to the cumbersome and rudimentary computations, we just present the result. The natural lower bound we arrive to the upper bound of the function $\kappa$ for the $p$ segments curve can be found to be
[…]
$\text{ as long as}\quad 0\notin\operatorname{co}(S).$
[…]
We propose the following methodology. Suppose we have $p\in\mathbb{N}$ segments with $p\geq 2$. Using equation (8) and given a maximum curvature $\kappa_{\max}>0$, we can compute for each pair of consecutive segments its respective [^3] The index $i$ ranges from the first pair to the last pair of consecutive segments. $\varepsilon_{i}>0$ such that, under the three-point-two-segment approximation, their curvatures are upper bounded. If $\varepsilon_{i}<\frac{1}{2}$ for all $i$, then equation (8) is exact, because only the two segments used for computing $\varepsilon_{i}$ contribute to the mollification at the junction point. In this case, take $\varepsilon=\max_{i}\varepsilon_{i}$, which is valid because the only dependence on $\varepsilon_{i}$ in the right-hand side of equation (8) is through $\frac{1}{\varepsilon_{i}}$; hence, $\varepsilon$ satisfies the bound for each pair of segments. If $\varepsilon_{i}>\frac{1}{2}$ for some $i$, then equation (8) becomes an approximation. In this case, one can either accept an admissible error or use $\varepsilon=\max_{i}\varepsilon_{i}$ as an initial condition for an optimization algorithm that seeks the minimum $\varepsilon>0$ that upper bounds the curvature. In either case, equation (8) is a powerful, computationally inexpensive tool that can be used to either compute an exact upper bound for the complete trajectory or reduce computation time in an optimization algorithm.
VI Numerical validations and real experiments
To demonstrate the effectiveness of our path generation approach, we present both numerical and experimental results for path following of a mollified path by a unicycle vehicle. Specifically, we employ the Singularity-Free Guiding Vector Fields (SF-GVF) path following algorithm [17, 14]. In brief, SF-GVF takes a parametric path $f\in C^{2}(\mathbb{R},\mathbb{R}^{n})$ as input and constructs a vector field $\chi\in C^{2}(\mathbb{R}^{n},\mathbb{R}^{n})$ whose flow traces the mollified path.
VI-A Numerical results
We are going to consider the so called “heart” function as our input path. Define the function
$t\in[0,2\pi)\to r(t)=2-2\sin(t)+\sin(t)\frac{\sqrt{|\cos(t)|}}{\sin(t)+1.4}.$
For $t\in[0,2\pi)$ let $f_{1}(t)=r(t)\cos(t)$ and $f_{2}(t)=r(t)\sin(t)$, and we call $f:=(f_{1},f_{2})$ the “heart” path. Note that the “heart” path is continuous but not differentiable; therefore, it cannot be used for the path following algorithm SF-GVF. We solve this issue by approximating the function using mollifiers. Let $\varphi$ the mollifier presented in Example 1 and let $\varepsilon_{1},\varepsilon_{2}>0$ be real numbers. We then mollify the “heart” path $F$ is defined as follows
$F=(F_{1},F_{2}):=(f_{1}*\varphi_{\varepsilon_{1}},f_{2}*\varphi_{\varepsilon_{2}}).$
Figure 8: Representation of the numerical simulation. In both pictures the original path $f$, is shown as a blue dotted line, while as a red dotted line the mollified trajectory $F$ for $\varepsilon_{1}=\varepsilon_{2}=0.4$ in the left picture, and for $\varepsilon_{1}=0.2$ and $\varepsilon_{2}=0.8$ in the right picture. The black solid line represents the (now smooth) flow generated by the guiding vector field according to [14] starting from an arbitrary initial position.
A numerical simulation of the vehicle under SF-GVF using $\varepsilon_{1}=\varepsilon_{2}=0.4$ for the “heart” path is shown in the left plot of Figure 8. The vehicle’s trajectory indicates convergence to the desired mollified path. Moreover, as $\varepsilon\to 0$, the trajectory approaches the original path more closely. Note that the mollified path lies inside the original path, as predicted by Theorem 6. In practical terms, this means that the “heart” function can now be used with SF-GVF, extending the applicability of this path following algorithm. Clearly, to improve convergence to the original path, we can reduce both $\varepsilon_{1}$ and $\varepsilon_{2}$, since by Theorem 1 we have uniform convergence on compact sets as $\varepsilon_{1},\varepsilon_{2}\to 0$. However, to demonstrate the flexibility of the approach, we also consider the case $\varepsilon_{1}=0.2$ and $\varepsilon_{2}=0.8$, whose simulation is presented in the right plot of Figure 8. Note that in the first component, the mollified curve is better adjusted to the original curve, while in the second component, the opposite occurs. This results from $\varepsilon_{2}$ being four times larger than $\varepsilon_{1}$. Indeed, the values of $\varepsilon_{1}$ and $\varepsilon_{2}$ can be constrained by the vehicle’s dynamics. This is a key advantage of the method: by simply adjusting these parameters, we can ensure that the vehicle follows the curve within its dynamic limits, thereby avoiding issues related to reconverging to the path. Moreover, numerical computations show that the length of the original path with respect to the $\ell_{1}$ norm is (in arbitrary units) $25.58$, while the mollified curve has length $23.16$ in the left case and $21.74$ in the right case of Figure 8, as predicted by Theorem 8. The same conclusions from Theorem 8 can be verified with respect to any other arbitrary $\ell_{p}$ norm.
Figure 9: Representation of the three dimensional numerical simulation. The original path $f$ is represented as a solid blue line, while the mollified trajectory $F=(F_{k})_{k=1}^{3}=(f_{k}*\varphi_{\varepsilon_{k}})_{k=1}^{3}$ is shown as a dashed red line, with $\varphi$ as in example 1 and $\varepsilon_{k}=1$ for $k\in\{1,2,3\}$. The black solid line represents the (now smooth) flow generated by the guiding vector field according to [14] starting from an arbitrary initial position.
Finally, for completeness, a numerical simulation of a 3D mollified path is presented in Figure 9. The notation used is identical to that in the previous numerical simulations. The original path $f$ is constructed via linear interpolation between a sequence of vertices/waypoints of a three-dimensional cube. As can be seen, $f$ is non-differentiable at these vertices. In contrast, the mollified function $F$ provides a smooth approximation that can be effectively employed in SF-GVF, as illustrated in Figure 9. Indeed, the flow of the guiding vector field converges to the mollified trajectory. Moreover, Theorem 8 can also be validated numerically. In this case, the length of the original path in the $\ell_{2}$ norm is (in arbitrary units) $7.38$, while the length of the mollified path is $5.48$. An important feature of this approach is its scalability: any trajectory of the form $g:\mathbb{R}\to\mathbb{R}^{n}$ can be mollified component-wise, producing a sufficiently smooth curve that can be further adapted to satisfy a variety of constraints.
VI-B Experimental results
Before presenting the experimental results, we introduce the software and hardware platforms used in the experiments. We also provide the necessary links to the developed software so any interested reader can replicate these experiments.
VI-B1 Software platform
We use Paparazzi UAV [18] as a development environment for programming autonomous vehicles. Paparazzi is a free and open-source hardware and software project designed to create a flexible autopilot system. Although it was initially developed for unmanned aerial vehicles (UAVs), researchers from various universities have extended its applications to other autonomous vehicles. Moreover, Paparazzi supports multiple hardware designs, making the interconnection between different modules accessible; the user simply needs to focus on the desired algorithm to implement, e.g., mollification of trajectories. Another important component of Paparazzi is its Ground Control Station (GCS), where vehicle positions are visualized and different sensor and internal variables can be monitored via telemetry. We have added the necessary code to define, display, and modify the mollification of straight lines on the fly. See Figure 10 for a screenshot of the Paparazzi GCS showing the mollification of a linear interpolation of points in $\mathbb{R}^{2}$.
Figure 10: Capture of an experiment using Paparazzi GCS. The original trajectory $f$ is shown in green created by linearly interpolating the points $L_{i}\in\mathbb{R}^{2}$, $i\in\{0,\dots,7\}$. The yellow curve represents the mollification of the original trajectory $F=(f_{1}*\varphi_{\varepsilon},f_{2}*\varphi_{\varepsilon})$ where $\varphi$ is as in Example 1 and $\varepsilon=0.5$. The orange line represents the trajectory described by the vehicle.
Using the Paparazzi GCS, the user can issue real-time high-level commands to the Rover’s autopilot via telemetry. Examples of such commands include path following instructions for desired waypoints, speed setpoints, and guidance and speed controller constants. The rest of the system runs on the vehicle’s onboard computer, enabling fully autonomous operation. It is important to note that waypoints can be adjusted to fit the trajectory into the desired area. By simply moving the control or endpoints, the user can shape the path as desired. Onboard software in a microcontroller recalculates the trajectory by mollifying it using the user-specified parameters whenever a point or parameter is changed. This provides considerable flexibility in experimental environments where positional uncertainties may arise.
VI-B2 Experimental platform
Figure 11: Rover vehicle and hardware used during the experiments.
The experimental hardware platform is presented in Figure 11 and can be modeled using the standard unicycle model. The rover is equipped with a Matek F765-Wing Autopilot as its central component. The autopilot consists of an STM32 microcontroller and integrated Inertial Measurement Units (IMUs) for inertial navigation, along with input/output ports for connecting various sensors including barometers, GNSS, compass, and radio receivers. We use the open-source Paparazzi UAV framework for programming and uploading code to the autopilot. Paparazzi supports the sensors required for outdoor autonomous operation: GNSS receivers and compass for accurate positioning and yaw attitude, radio receivers as a fail-safe for manual control during algorithmic failures, and telemetry radios enabling ground reception of vehicle states and transmission of commands to the vehicle. Our implementation uses the Matek M10Q-5883 GNSS and compass module, a Futaba 7008SB radio receiver, and a Zigbee Xbee telemetry radio.
VI-B3 Experimental results
Experimental data extracted from Paparazzi logs is shown in Figure 12. We created a continuous but non-differentiable path by linearly interpolating points in $\mathbb{R}^{2}$ as it is done in Section V. As noted above, any Paparazzi user can create such trajectories by simply moving points in the ground control station before or during the experiment in real time. The experiment uses a curve similar to that in Figure 10.
The experimental objective is as follows: Given a desired non-differentiable path $f$, mollify it with parameter $\varepsilon>0$ to obtain $F_{\varepsilon}=(f*\varphi_{\varepsilon})$, where $\varepsilon$ is a function of vehicle speed limited by maximum allowed curvature. To demonstrate potential applications, we use a linear relationship between maximum allowed curvature and speed. For speed $v>0$, with $R_{min}$, $R_{max}$, and $v_{max}$ denoting minimum radius, maximum radius, and maximum speed respectively, the allowed radius of curvature is $R(v):=R_{min}+\frac{v}{v_{max}}(R_{max}-R_{min})$. Note this is merely illustrative; curvature and speed need not be linearly related in practice.
We adapt the curve to vehicle dynamics as follows: knowing the speed at each segment midpoint and assuming it remains constant until the next segment midpoint, we compute maximum allowed curvature using the linear relationship, then obtain the minimum allowed $\varepsilon>0$ from equation (8). This process applied to each segment pair creates a family of curves for different $\varepsilon>0$ values. With six segments, we obtain five mollified curves (one per corner) plus one for the default $\varepsilon>0$ value. As before, $f:\mathbb{R}\to\mathbb{R}^{2}$ denotes the original path, $\varphi$ is the mollifier from Example 1, $F_{\varepsilon}=(f*\varphi_{\varepsilon})$ is the speed-dependent mollified path, and $r:[0,\infty)\to\mathbb{R}^{2}$ denotes vehicle position as a function of time, with $r(t;r_{0})$ indicating initial position $r_{0}\in\mathbb{R}^{2}$.
Figure 12 shows the results. The family of six mollified paths adapts well to vehicle dynamics, approaching the original path when curvature constraints allow (e.g., segments four and five) while strongly constraining the curve when necessary (e.g., the last two segments). The theoretical results hold: mollified curves lie within the convex hull of the original path (Theorem 6), and parameters are sufficiently small that Propositions 2 and 5 apply. These experiments validate our approach, demonstrating both theoretical solution of Problem 1 and practical viability on real and affordable hardware.
Figure 12: Experiments with the real rover vehicle. Left column figures represent the original trajectory in black, the mollified family trajectories as red dashed lines, the complete trajectory of the vehicle as a dashed blue line, its initial point as a blue cross, and the position of the vehicle for its corresponding $\varepsilon_{i}>0$ in a thick solid blue line. Right column figures represent in a similar fashion, the error between the rover’s position and the path in blue, and the GNSS accuracy in green.
VII Conclusions
In this work, we addressed the problem of efficient path generation to make non-suitable curves, such as linear interpolations from waypoint collections, suitable for path following and trajectory tracking algorithms via mollification. The mollification can be adjusted so that the mollified trajectory approximates the original trajectory arbitrarily closely on compact sets while being completely smooth. Additionally, properties such as convexity, concavity, monotonicity, and quasiconvexity are preserved under mollification, with local versions also preserved for sufficiently small mollification parameters.
We validated the approach through numerical simulations using Singularity-Free Guiding Vector Fields as a path following algorithm, applying mollification to the “heart” path and a 3D trajectory while examining the effects of different parameter values. Finally, experiments on rovers demonstrated the viability of the approach and the ability to tune mollified paths to match vehicle dynamics. While the original trajectory may not be physically realizable, the mollified trajectory with appropriate parameters can be successfully followed. This confirms that our results have both theoretical significance and practical value for autonomous vehicles, industrial robotics, and any engineering application requiring fast and rigorous function approximation.

### v2

@@
-Efficient Generation of Smooth Paths with Curvature Guarantees by Mollification
+Efficient Path Generation with Curvature Guarantees by Mollification
 Abstract
-Most path following and trajectory tracking algorithms in mobile robotics require the desired path or trajectory to be defined by at least twice continuously differentiable functions to guarantee key properties such as global convergence, especially for nonholonomic robots like unicycles with speed constraints.
-Consequently, these algorithms typically exclude continuous but non-differentiable paths, such as piecewise functions.
-Despite this exclusion, such paths provide convenient high-level inputs for describing robot missions or behavior.
-While techniques such as spline interpolation or optimization-based methods are commonly used to smooth non-differentiable paths or create feasible ones from sequences of waypoints, they either can produce unnecessarily complex trajectories or are computationally expensive.
+Path generation, the process of converting high-level mission specifications, such as sequences of waypoints from a path planner, into smooth, executable paths, is a fundamental challenge in mobile robotics.
+Most path following and trajectory tracking algorithms require the desired path to be defined by at least twice continuously differentiable functions to guarantee key properties such as global convergence, especially for nonholonomic robots like unicycles with speed constraints.
+Consequently, path generation methods must bridge the gap between convenient but non-differentiable planning outputs, such as piecewise linear segments, and the differentiability requirements imposed by downstream control algorithms.
+While techniques such as spline interpolation or optimization-based methods are commonly used to smooth non-differentiable paths or create feasible ones from sequences of waypoints, they either produce unnecessarily complex trajectories or are computationally expensive.
 In this work, we present a method to regularize non-differentiable functions and generate feasible paths through mollification.
@@
 Additionally, we provide a systematic method for bounding the curvature of generated paths, which we demonstrate by applying it to paths resulting from linking a sequence of waypoints with segments.
-The proposed approach is computationally efficient, enabling real-time implementation on microcontrollers and compatibility with standard trajectory tracking and path following algorithms.
-I Introduction
-Path generation transforms a set of points, e.g., a finite collection of waypoints, into a smooth, continuous curve that satisfies geometric properties such as continuity, bounded curvature, and differentiability, to respect vehicle kinematic constraints, ultimately transforming a sparse set of discrete locations into a followable geometric path.
-Spline-based methods, including cubic B-splines [1, 2] and Bézier curves [3], ensure $C^{2}$ continuity and bounded curvature but may produce high curvature and cumbersome paths when the spacing of waypoints is irregular [4].
+The proposed approach is analytically shown to be computationally more efficient than standard interpolation methods, enabling real-time implementation on microcontrollers, while remaining compatible with standard trajectory tracking and path following algorithms.
+1 Introduction
+In robotics and control, motion execution is often decomposed into a hierarchy of interrelated problems.
+Path planning determines a feasible route through the robot’s configuration space—typically as a sparse sequence of waypoints—using methods ranging from graph-search algorithms like A* to sampling-based approaches such as RRT [10].
+Path generation then transforms this abstract plan into a smooth, continuous geometric path or time-parameterized trajectory, often using Spline-based methods, including cubic B-splines [9, 1] and Bézier curves [16].
+These methods ensure $C^{2}$ continuity and bounded curvature but may produce high curvature and cumbersome paths when the spacing of waypoints is irregular [11].
 Gaussian process regression [13] and kernel smoothing [5] generate $C^{\infty}$-smooth approximations, trading exact waypoint passage for improved smoothness at the cost of $O(n^{3})$ computational complexity and potential numerical instability.
 Optimization-based approaches [12, 14] incorporate explicit curvature bounds through convex formulations, enabling real-time computation but requiring careful parameter tuning [7].
-Sampling-based methods, like RRT* [10] with post-smoothing offer probabilistic completeness but may yield suboptimal curvature or other metrics.
-In summary, it is not surprising that the chosen method balances curvature control, computational complexity, and application requirements for interpolation versus approximation.
+However, critically, path generation does not operate in isolation: its output must meet the requirements imposed by the downstream control layer, which typically takes one of two forms.
+In path following, the generated path must be geometrically smooth and continuously differentiable, so that a well-defined tangent and curvature exist at every point for spatial error computation [17].
+In trajectory tracking, the demands are more stringent: the time parameterization must additionally be consistent with the robot’s dynamic capabilities, as a reference trajectory that demands accelerations or velocities beyond the system’s limits cannot be reliably tracked regardless of the controller’s design [15].
+In both cases, path generation must internalize the robot’s kinematic and dynamic constraints, effectively acting as the bridge that makes a planned route physically executable and controllable.
 In this paper, we propose an inexpensive path generation method based on mollifying non-differentiable paths, e.g., piecewise functions, using mollifier functions.
@@
 However, to the best of the authors’ knowledge, mollifiers have not been applied to path generation problems, despite their natural fit for robotics applications.
-For example, given a finite collection of ordered waypoints, it produces a path that can be arbitrarily close, in the Euclidean sense, to the segments connecting the waypoints sequentially.
+Mollification offers a principled approach to path generation.
+For example, given a finite collection of ordered waypoints, it produces a smooth curve that approximates the piecewise-linear route to an arbitrary degree of Euclidean closeness.
 For this specific generated path, our method is well-suited for unicycle-like robots that travel at constant speed with heading-rate constraints; i.e., we can guarantee a maximum curvature for the generated path at the mild cost of slightly deviating from these connecting segments—an unavoidable trade-off, since the union of segments produces a non-differentiable path.
@@
 We also show how our method can generate 3D paths and guarantee several properties for generic sets of input points, including convexity preservation, enclosure of the generated path, and bounded maximum length when the input is also closed.
-The article is organized as follows, Section III introduces the mollifier functions and the problem formulation.
-Section IV explores which geometrical and analytical aspects of the original function are affected by mollification.
+The article is organized as follows, Section 3 introduces the mollifier functions and the requirements for the path generation problem.
+Section 4 explores which geometrical and analytical aspects of the original function are affected by mollification.
 Convexity, concavity and quasiconvexity are treated, as well as the length of the path and the curvature.
-Section VI validates the theoretical findings numerically and experimentally.
-We end the article with some conclusions in Section VII.
-II Notation
+Section 6 validates the theoretical findings numerically and experimentally.
+We end the article with some conclusions in Section 7.
+2 Notation
 In this paper we use Lebesgue integration rather than Riemann integration to overcome some limitations with sets of measure zero and interchanging integral and limits.
-Therefore, we denote the Lebesgue measure in one dimension as $\lambda$, i.e., $\lambda:=\lambda_{1}$, and as $\lambda_{n}$ the $n$’th dimensional Lebesgue measure, and we consider the measure space $(X,\mathcal{B}(X),\lambda_{n})$ with $X\subseteq\mathbb{R}^{n}$, and $\mathcal{B}(X)$ is the Borel sigma-algebra of $X$; hence, when we say a function is measurable we are actually saying that it is Borel measurable.
+Therefore, we denote the Lebesgue measure in one dimension as $\lambda$, i.e., $\lambda:=\lambda_{1}$, and as $\lambda_{n}$ the $n$’th dimensional Lebesgue measure, and we consider that it is equipped with the Borel sigma algebra.
 Consequently, all integrals in this paper must be thought as the Lebesgue integral with respect to the Lebesgue measure, even if the notation $\lambda$ is sometimes omitted in the integral.
-For $p\in[1,\infty)$, we say that a measurable function $f$ is locally $p$-integrable, denoted $f\in L^{p}_{\mathrm{loc}}(X)$, if for each $x\in X$ there exists a neighborhood $U$ of $x$ such that $f|_{U}\in L^{p}(U)$.
-We denote by $\operatorname{id}:X\to X$ the identity function defined as $\operatorname{id}(x)=x$. Then, the indicator function of a set $A$ is defined as
+For $p\in[1,\infty)$, we say that a measurable function $f$ is locally $p$-integrable, denoted $f\in L^{p}_{\mathrm{loc}}(X)$, if it is $p$-integrable in each compact subset of $X$.
+We denote by $\operatorname{id}:X\to X$ the identity function defined as $\operatorname{id}(x)=x$.
+Then, the indicator function of a set $A$ is defined as
 $\operatorname{ind}_{A}(x)=\begin{cases}1&x\in A,\\ 0&x\notin A\end{cases}.$
@@
 For a set $A$ we denote its closure as $\overline{A}$ and the support of a real-valued function is defined as $\operatorname{supp}f:=\overline{\{x\in\operatorname{dom}f\mid f(x)\neq 0\}}$, where $\operatorname{dom}f$ is the domain of $f$.
-Finally, a (parametric) path in $\mathbb{R}^{n}$ is a function $f:X\subset\mathbb{R}\to\mathbb{R}^{n}$.
+Finally, a (parametric) path in $\mathbb{R}^{n}$ is a measurable function $f:X\subset\mathbb{R}\to\mathbb{R}^{n}$.
 Writing $f=(f_{1},\dots,f_{n})$, we call $f_{i}$ the $i$’th component of the path for $i\in\{1,\dots,n\}$.
-III Path generation by mollification
-III-A Path generation requirements
+3 Path generation by mollification
+3.1 Path generation requirements
 This paper seeks an alternative approach to interpolation and optimization methods for the generation of paths from high-level inputs—such as waypoint sequences—that is computationally efficient, conceptually simple, and has a transparent physical interpretation.
@@
 Problem 1 (Path generation problem)
-Let $f:\mathbb{R}\to\mathbb{R}^{n}$ be a measurable path and $\{\varepsilon_{i}\}_{i=1}^{n}$ be a collection of positive real numbers.
-Find a new path $T_{\varepsilon}(f):\mathbb{R}\to\mathbb{R}^{n}$ where $T_{\varepsilon}(f)=:\{T_{\varepsilon_{i}}(f)\}_{i=1}^{n}$ and $T_{\varepsilon_{i}}$ is a functional that acts on each component of $f$ such that:
+Let $f:\mathbb{R}\to\mathbb{R}^{n}$ be a parametric path and $\{\varepsilon_{i}\}_{i=1}^{n}$ be a collection of positive real numbers.
+Find a new path $T_{\varepsilon}(f):\mathbb{R}\to\mathbb{R}^{n}$ where $T_{\varepsilon}(f)=:\{T_{\varepsilon_{i}}(f_{i})\}_{i=1}^{n}$ and $T_{\varepsilon_{i}}$ is a functional that acts on each component of $f$ such that:
   1. 1.
-    $T_{\varepsilon_{i}}(f_{i})\to f_{i}$ as $\varepsilon_{i}\to 0$ in some sense of convergence.
+    Each component of the parametric path can be made arbitrary close to the original path, that is, $T_{\varepsilon_{i}}(f_{i})\to f_{i}$ as $\varepsilon_{i}\to 0$ in some sense of convergence.
   2. 2.
-    For $p\in\mathbb{N}$ with $p\geq 2$, $T_{\varepsilon_{i}}(f_{i})\in C^{p}(\mathbb{R},\mathbb{R})$ for $i\in\{1,\dots,n\}$.
+    It provides enough smoothness, that is, for $p\in\mathbb{N}$ with $p\geq 2$, $T_{\varepsilon_{i}}(f_{i})\in C^{p}(\mathbb{R},\mathbb{R})$ for $i\in\{1,\dots,n\}$.
   3. 3.
@@
 Finally, the third requirement enables real-time path generation with low-cost hardware.
-III-B Mollifiers for the path generation
+3.2 Mollifiers for the path generation
 The solution to Problem 1 can be obtained by taking a weighted average of the points along the parametric path $f$ through convolution with a certain type of function known as a mollifier [4].
@@
 Now, let us consider a set of mollifiers $\{\varphi_{\varepsilon_{i}}:\mathbb{R}\to\mathbb{R}\}_{i=1}^{n}$, and define the generated path $F=(F_{1},\dots,F_{n})$ with $F_{i}:=f_{i}*\varphi_{\varepsilon_{i}}$ for all $i\in\{1,\dots,n\}$, and let us denote by $\varphi_{\varepsilon}^{(n)}$ the $n$-th derivative of the function $\varphi_{\varepsilon}$.
-Theorem 1 ([11, Appendix C, Theorem 7])
+Theorem 1 ([4, Appendix C, Theorem 7])
 Let $\varepsilon>0$, $f\in L_{loc}^{p}(\mathbb{R})$ with $p\in[1,\infty]$, and $\varphi$ be a mollifier; then the following three statements hold:
@@
 Indeed, the second requirement is satisfied straightforwardly.
-Regarding the computational efficiency from the third requirement, the computation of $F$ is straightforward.
+Regarding the computational efficiency from the third requirement, the computation of $F$ is also straightforward.
 Indeed, because of the compact and symmetric support of the mollifier due to Assumption 1, we have that $F_{\varepsilon}(x)=\int_{[-\varepsilon,\varepsilon]}f(x-t)\varphi_{\varepsilon}(t)\,\mathrm{d}\lambda(t)$ is an inexpensive numerical operation, and note the compact integration interval.
@@
     If $f$ is continuous then $F_{\varepsilon}\to f$ as $\varepsilon\to 0$ uniformly on compact subsets of $\mathbb{R}$.
-Proof:
+Proof 3.3.
 For the first statement, suppose $f$ is uniformly continuous.
@@
 Thus, the same arguments as above can be followed noting that in this case $\delta$ depends on the compact set $K$.
-∎
-IV Key properties of the generated path
+4 Key properties of the generated path
 While Theorems 1 and 2 solve Problem 1, they do not provide further details about the properties of the resulting path.
@@
 How is the output path enclosed, and what is its length when the input is a closed path?
-We defer the analysis of the output path’s curvature to Section V, where we provide a detailed curvature analysis for the case when the input is a sequence of 2D or 3D waypoints.
+We defer the analysis of the output path’s curvature to Section 5, where we provide a detailed curvature analysis for the case when the input is a sequence of 2D or 3D waypoints.
 This curvature analysis serves as a systematic methodology that can be applied to other types of input paths.
 […]
-IV-A Convexity properties
+4.1 Convexity properties
 Convexity and local convexity are properties of great interest in the study of trajectory shapes.
@@
 We now present several propositions and counterexamples addressing these questions.
-Proposition 1 (Convexity and mollification)
+Proposition 4.4 (Convexity and mollification).
 Let $f\in L^{1}_{loc}(\mathbb{R})$ be convex and $\varphi$ a nonnegative mollifier as in Definition 2; then $F_{\varepsilon}:=\varphi_{\varepsilon}*f$ is convex for any $\varepsilon>0$.
-Proof:
+Proof 4.5.
 Let $x,y\in\mathbb{R}$ and $\gamma\in[0,1]$. Since $\varphi_{\varepsilon}\geq 0$ we have that
@@
 We now present a proposition and a counterexample.
-Proposition 2 (Local convexity and mollification)
+Proposition 4.6 (Local convexity and mollification).
 Let $f\in L^{1}_{loc}(\mathbb{R})$ be a function that is convex in some set $(a,b)\subseteq\mathbb{R}$ with $\infty\leq a<b\leq\infty$.
@@
 Then, for each $x,y\in(a,b)$ with $x<y$ there exists a $\delta=\delta(x,y)>0$ such that for all $\varepsilon\in(0,\delta)$ the function $F_{\varepsilon}:=f*\varphi_{\varepsilon}$ is convex in the set $(x,y)$.
-Proof:
+Proof 4.7.
 Let $x,y\in(a,b)$.
@@
 Since $V$ is convex we have that $\gamma(\xi-t)+(1-\gamma)(\zeta-t)\in V$ for all $t\in(-\varepsilon,\varepsilon)$ and $\gamma\in[0,1]$.
-Given that $f$ is convex in $(a,b)$ it is also convex in $V\subset(a,b)$, and noting $\varphi_{\varepsilon}\geq 0$ we can follow the steps of the proof of proposition 1 to reach
+Given that $f$ is convex in $(a,b)$ it is also convex in $V\subset(a,b)$, and noting $\varphi_{\varepsilon}\geq 0$ we can follow the steps of the proof of proposition 4.4 to reach
 $F_{\varepsilon}(\gamma\xi+(1-\gamma)\zeta)$ $\leq\gamma F_{\varepsilon}(\xi)+(1-\gamma)F_{\varepsilon}(\zeta).$
-Because $\xi,\zeta\in V$ and $\gamma\in[0,1]$ were arbitrary the proposition follows. ∎
-Example 2
-It is natural to ask whether Proposition 2 holds for any $\varepsilon>0$.
+Because $\xi,\zeta\in V$ and $\gamma\in[0,1]$ were arbitrary the proposition follows.
+Example 4.8.
+It is natural to ask whether Proposition 4.6 holds for any $\varepsilon>0$.
 That is, if $f$ is convex on an open set, is $F_{\varepsilon}$ also convex on that open set independently of $\varepsilon$ and the choice of non-negative mollifier?
@@
 Nevertheless, for $\varepsilon=0.45$, we have that $F_{\varepsilon}$ is convex on a neighborhood contained in (0.5,0.5)(-0.5,0.5).
-Thus, we can set $\delta=0.45$ in Proposition 2, and for any $\varepsilon<\delta$, $F_{\varepsilon}$ is convex in (0.5,0.5)(-0.5,0.5).
-Remark 1
-Note that Propositions 1 and 2 also hold when $f$ is concave instead of convex, with concavity replacing convexity throughout.
+Thus, we can set $\delta=0.45$ in Proposition 4.6, and for any $\varepsilon<\delta$, $F_{\varepsilon}$ is convex in (0.5,0.5)(-0.5,0.5).
+Remark 4.9.
+Note that Propositions 4.4 and 4.6 also hold when $f$ is concave instead of convex, with concavity replacing convexity throughout.
 Before the following result, we need to prove that affine maps are invariant under mollification.
-Proposition 3 (Affine functions and mollification)
+Proposition 4.10 (Affine functions and mollification).
 Let $\varphi_{\varepsilon}$ with $\varepsilon>0$ be a mollifier with symmetric support around the origin, and let $a,b\in\mathbb{R}$; then $\varphi_{\varepsilon}*(a\operatorname{id}+b)=a\operatorname{id}+b$.
-Proof:
+Proof 4.11.
 Let $x\in\mathbb{R}$ and $\varepsilon>0$. Then
@@
 $=ax+b-a\int_{(-\varepsilon,\varepsilon)}y\varphi_{\varepsilon}(y)\,\mathrm{d}\lambda(y).$
-The result follows noting that $y\in\mathbb{R}\mapsto y\varphi_{\varepsilon}(y)$ is an odd function that is integrated over a symmetric interval. ∎
+The result follows noting that $y\in\mathbb{R}\mapsto y\varphi_{\varepsilon}(y)$ is an odd function that is integrated over a symmetric interval.
 We also need the Jensen’s inequality.
-Theorem 3 (Jensen’s inequality)
-[15, Theorem 1.6.2] Let $\varphi$ be a non-negative measurable function such that $\int_{\mathbb{R}}\varphi\mathrm{d}\lambda=1$, $g$ be any measurable function and $f$ be a convex function such that $\operatorname{dom}f\supset\operatorname{img}g$; then
+Theorem 4.12 (Jensen’s inequality).
+[3, Theorem 1.6.2] Let $\varphi$ be a non-negative measurable function such that $\int_{\mathbb{R}}\varphi\mathrm{d}\lambda=1$, $g$ be any measurable function and $f$ be a convex function such that $\operatorname{dom}f\supset\operatorname{img}g$; then
 $f\left(\int_{\mathbb{R}}g(x)\varphi(x)\mathrm{d}\lambda(x)\right)\leq\int_{\mathbb{R}}(f\circ g)(x)\varphi(x)\mathrm{d}\lambda(x).$
@@
 Figure 3: Mollification of the function $f=|\operatorname{id}|$ with the mollifier $\varphi_{\varepsilon}$ as in Example 1 for different values of $\varepsilon$.
-Note that the function $f$ is convex, implying $F_{\varepsilon}$ is convex as demonstrated in proposition 1 and it is above the graph as shown in Proposition 4.
+Note that the function $f$ is convex, implying $F_{\varepsilon}$ is convex as demonstrated in proposition 4.4 and it is above the graph as shown in Proposition 4.13.
 Now, we are ready to show that if $f$ is convex and $\varphi_{\varepsilon}$ is a non-negative even mollifier, then $\varphi_{\varepsilon}*f\geq f$ pointwise.
-For example, combined with Proposition 3, the following result implies that if our trajectory resembles an inverted tent, then the mollified function will also resemble a smoothed inverted tent but with the mollified function lying above the original path; see Figure 3 for the illustration.
-Proposition 4 (If $f$ is convex then $F_{\varepsilon}$ is above $f$)
+For example, combined with Proposition 4.10, the following result implies that if our trajectory resembles an inverted tent, then the mollified function will also resemble a smoothed inverted tent but with the mollified function lying above the original path; see Figure 3 for the illustration.
+Proposition 4.13 (If $f$ is convex then $F_{\varepsilon}$ is above $f$).
 Let $\varphi_{\varepsilon}$ with $\varepsilon>0$ satisfying Assumption 1, $f$ be a convex function and define $F_{\varepsilon}:=\varphi_{\varepsilon}*f$. Then $F_{\varepsilon}\geq f$.
-Proof:
+Proof 4.14.
 Note that $\varphi_{\varepsilon}$ satisfies the conditions in Jensen’s inequality and that $f$ is convex.
@@
 $=f((\varphi_{\varepsilon}*\operatorname{id})(x)).$
-The result now follows from Proposition 3 since $x$ is arbitrary. ∎
-Remark 2
+The result now follows from Proposition 4.10 since $x$ is arbitrary.
+Remark 4.15.
 Note that a similar result also holds if $f$ is concave, where clearly $F_{\varepsilon}$ would be below $f$ in that case.
-The following result resembles Proposition 2, i.e., in order for the mollified function to be above the original function the parameter $\varepsilon$ has to be sufficiently small.
-Proposition 5
+The following result resembles Proposition 4.6.
+Proposition 4.16 (Local convexity and $F_{\varepsilon}\geq f$).
 Let $\varphi$ be even and satisfying Assumption 1, and $f:\mathbb{R}\to\mathbb{R}$ be a function that is convex in $(a,b)$.
 Then, for each $x,y\in(a,b)$ with $x<y$ there exists a $\delta=\delta(x,y)>0$ such that for all $\varepsilon\in(0,\delta)$ we have that $F_{\varepsilon}=\varphi_{\varepsilon}*f\geq f$ on $(x,y)$.
-Proof:
+Proof 4.17.
 Let $x,y\in(a,b)$.
@@
 $=f((\varphi_{\varepsilon}*\operatorname{id})(\xi))=f(\xi),$
-where the last equality comes from Proposition 3. Since $\xi\in(x,y)$ was arbitrary the result follows. ∎
-Remark 3
-Note that as it can be seen in Example 2, Proposition 5 does not hold for any $\varepsilon>0$.
-Indeed, also note that if the function is locally concave then we reach a similar results where the original function will be above the mollified function.
-IV-B Quasiconvexity under mollification
+where the last equality comes from Proposition 4.10. Since $\xi\in(x,y)$ was arbitrary the result follows.
+Remark 4.18.
+Note that as it can be seen in Example 4.8, Proposition 4.16 does not hold for any $\varepsilon>0$.
+Indeed, also note that if the function is locally concave then we reach a similar result where the original function will be above the mollified function.
+4.2 Quasiconvexity under mollification
 Any convex function is quasiconvex, but the converse does not hold in general.
@@
 From an engineering standpoint, this is particularly useful for estimating the shape of the mollified curve.
-Definition 3 (Quasiconvex function)
+Definition 4.19 (Quasiconvex function).
 Let $S$ be a non empty convex set and $f:S\to\mathbb{R}$ be a real-valued function. We say that $f$ is quasiconvex if for all $\alpha\in\mathbb{R}$ the set
@@
 We need the following lemma before proving that quasiconvexity is preserved under mollification.
-Lemma 4
+Lemma 4.20.
 Let $f:\mathbb{R}\to\mathbb{R}$ be a quasiconvex function, $a,b\in\mathbb{R}$, and define $g:=a\operatorname{id}+b$; then, the following statements hold:
@@
     $f\circ g$ is quasiconvex.
-Proof:
+Proof 4.21.
 First we prove that $f$ is measurable.
@@
 $=\alpha.$
-That is, $\gamma x+(1-\gamma)y\in S_{\alpha}$. Therefore $S_{\alpha}$ is a convex set, proving $f\circ g$ is quasiconvex. ∎
+That is, $\gamma x+(1-\gamma)y\in S_{\alpha}$. Therefore $S_{\alpha}$ is a convex set, proving $f\circ g$ is quasiconvex.
 Having presented the main properties of quasiconvex functions, let us consider a representative quasiconvex case, the monotonic function.
 Interesting enough, the mollified function of a monotonic path, e.g., a staircase-like sequence of steps, will also be monotonic.
-Proposition 6 (Monotonicity and mollification)
+Proposition 4.22 (Monotonicity and mollification).
 Let $f\in L^{1}_{loc}(\mathbb{R})$ be monotone increasing (resp.
@@
 decreasing).
-Proof:
+Proof 4.23.
 Suppose $f$ is monotone increasing. Let $x,y\in\mathbb{R}$ with $x>y$. Then for any $t\in\mathbb{R}$ we have xt>ytx-t>y-t, thus $f(x-t)\geq f(y-t)$.
 $F_{\varepsilon}(x)-F_{\varepsilon}(y)$ $=\int_{\mathbb{R}}[f(x-t)-f(y-t)]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)\geq 0,$
-since $\varphi_{\varepsilon}$ is positive. The proof is identical for monotone decreasing functions. ∎
+since $\varphi_{\varepsilon}$ is positive. The proof is identical for monotone decreasing functions.
 Figure 4: Mollification of stair-step function created using shifted Heaviside step functions.
-The blue solid line represents the (discontinuous) stair-step function $f$, while the red solid line represents its mollification using the mollifier of equation (1) and $\varepsilon=0.5$.
-Due to proposition 6, $F_{\varepsilon}$ cannot present overshoots or oscillations.
+The blue solid line represents the (discontinuous) stair-step function $f$, while the red solid line represents its mollification using the mollifier (1) and $\varepsilon=0.5$.
+Due to Proposition 4.22, $F_{\varepsilon}$ cannot present overshoots or oscillations.
 Suppose the desired trajectory is the Heaviside step function $h:\mathbb{R}\to\mathbb{R}$ defined by $h(t)=\operatorname{ind}_{(0,\infty)}(t)$.
@@
 We finally show that quasiconvexity is preserved under mollification with a nonnegative mollifier.
-Theorem 5
-Suppose $f:\mathbb{R}\to\mathbb{R}$ is a quasiconvex function and let $\varphi$ be a mollifier satisfying 1; then, for all $\varepsilon>0$ the function $F_{\varepsilon}:=f*\varphi_{\varepsilon}$ is quasiconvex.
-Proof:
+Theorem 4.24.
+Suppose $f:\mathbb{R}\to\mathbb{R}$ is a quasiconvex function and let $\varphi$ be a mollifier satisfying Assumption 1; then, for all $\varepsilon>0$ the function $F_{\varepsilon}:=f*\varphi_{\varepsilon}$ is quasiconvex.
+Proof 4.25.
 We are going to proceed by contradiction.
@@
 First, we claim that the set $V_{1}:=\{t\in(-\varepsilon,\varepsilon)\mid f(\gamma x+(1-\gamma)y-t)-f(x-t)>0\}$ is measurable and it has positive measure.
-By Lemma 4 we know that the composition of a quasiconvex function with an affine mapping is quasiconvex, hence measurable.
+By Lemma 4.20 we know that the composition of a quasiconvex function with an affine mapping is quasiconvex, hence measurable.
 Since the linear combination of two measurable function results in a measurable function, the set $V_{1}$ is measurable.
-Thus, if $\lambda(V_{1})$ were zero, then it would negligible in integration, which would imply that
+Thus, if $\lambda(V_{1})$ were zero, then it would be negligible in integration, which would imply that
 $0$ $<\int_{(-\varepsilon,\varepsilon)}\left[f(\gamma x+(1-\gamma)y-t)-f(x-t)\right]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)$
@@
 but $(-\varepsilon,\varepsilon)\setminus V_{1}=\{t\in(-\varepsilon,\varepsilon)\mid t\notin V_{1}\}=\{t\in(-\varepsilon,\varepsilon)\mid f(\gamma x+(1-\gamma)y-t)-f(x-t)\leq 0\}$ and $\varphi$ is nonnegative, thus
-$\int_{(-\varepsilon,\varepsilon)\setminus V_{1}}\left[f(\gamma x+(1-\gamma)y-t)-f(x-t)\right]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)\leq 0.$
+$\int_{(-\varepsilon,\varepsilon)\setminus V_{1}}\left[f(\gamma x+(1-\gamma)y-t)-f(x-t)\right]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)\leq 0,$
 which leads to the first contradiction. Therefore $\lambda(V_{1})>0\Longrightarrow V_{1}\neq\emptyset$. Moreover, since $f$ is quasiconvex we have that for any $t\in V_{1}$
@@
 $\leq\max\{f(y-t_{1}),f(y-t_{2})\}$
-However, $f(y-t_{2})<f(x-t_{2})$ by equation (4), thus
+However, $f(y-t_{2})<f(x-t_{2})$ because of (4), thus
 $f(y-\beta t_{1}-(1-\beta)t_{2})<\max\{f(y-t_{1}),f(x-t_{2})\},$
 $\quad\forall\beta\in[0,1].$
-Using the same approach and considering equation (3) we have that
+Using the same approach and considering (3) we have that
 $f(x-\beta t_{1}-(1-\beta)t_{2})<\max\{f(y-t_{1}),f(x-t_{2})\},$
@@
 That is, $F_{\varepsilon}$ is quasiconvex for any $\varepsilon>0$.
-∎
-Remark 4
+Remark 4.26.
 The converse is not in general true.
@@
 Moreover, it is clear that if $F_{\varepsilon}$ is quasiconvex for all $\varepsilon>0$ then $f$ is quasiconvex since the pointwise limit of a family of quasiconvex functions can be shown to be quasiconvex function as well.
-IV-C Closed paths
+4.3 Enclosure and length of paths
 So far, we have worked with real-valued functions because we represent desired paths or trajectories as parametric functions, i.e., functions $f:\mathbb{R}\to\mathbb{R}^{n}$.
-Nevertheless, as previously noted, our results can be extended to functions $\phi:\mathbb{R}^{n}\to\mathbb{R}$, for which convexity-like properties can also be defined.
-This extension is of practical interest, since in some situations the desired path is a set of the form
-$\mathcal{P}=\{x\in\mathbb{R}^{n}\mid\phi_{i}(x)=0,\quad i\in\{1,\dots,n-1\}\},$
-for which previous results applying to each $\phi_{i}$.
-Nevertheless, there are situations in which the path arises from a parametric curve with compact domain, say $f:[a,b]\to\mathbb{R}^{n}$ where $\infty<a<b<\infty$, and it is preferable to work with $f$ as a whole rather than component by component.
-Since we lose some order-like properties in $\mathbb{R}^{n}$ relative to $\mathbb{R}$, we cannot directly extend convexity-like properties [^1] We refer to pure convexity properties..
-Nonetheless, we still want to characterize (in advance) how mollification affects the complete function $f$.
+Nonetheless, we still want to characterize (in advance) how mollification affects the complete function $f$, that is, treat it as a whole.
 We now address the following question: given the original trajectory, does there exist a subset $U$ of $\mathbb{R}^{n}$ such that the mollified trajectory is contained in $U$ for any value of its parameter?
-The answer is affirmative, with $U$ being the convex hull of $f([a,b])$.
-Definition 4
+The answer is affirmative, with $U$ being the convex hull of $f(\mathbb{R})$.
+Definition 4.27.
 Let $A\subset\mathbb{R}^{n}$ be a set. Its convex hull, denoted as $\operatorname{co}(A)$ is defined as the smallest convex set that contains $A$, that is, $A\subset\operatorname{co}(A)$. $\square$
 We first present a result when $\operatorname{dom}f=\mathbb{R}$.
-Theorem 6
+Theorem 4.28.
 Let $f:\mathbb{R}\to\mathbb{R}^{n}$ be a measurable function and $\varphi$ be a nonnegative mollifier. Define for $t\in\mathbb{R}$ and $\varepsilon>0$
-$F_{\varepsilon}(t)=(f*\varphi_{\varepsilon})(t)=\left((f_{1}*\varphi_{\varepsilon})(t),\dots,(f_{n}*\varphi_{\varepsilon})(t)\right).$
+$F_{\varepsilon}(t):=(f*\varphi_{\varepsilon})(t)=\left((f_{1}*\varphi_{\varepsilon})(t),\dots,(f_{n}*\varphi_{\varepsilon})(t)\right).$
 Then, given $\varepsilon>0$, we have that
-$\left\{F_{\varepsilon}(t)\mid t\in\operatorname{dom}f\right\}\subset\operatorname{co}\left\{f(t)\mid t\in\operatorname{dom}f\right\}.$
-Proof:
-Let $U=\operatorname{co}\{f(t)\mid t\in\operatorname{dom}f\}$. Define the extended real valued function $I_{U}:\mathbb{R}^{n}\to\mathbb{R}\cup\{-\infty,\infty\}$ as
+$\left\{F_{\varepsilon}(t)\mid t\in\mathbb{R}\right\}\subset\operatorname{co}\left\{f(t)\mid t\in\mathbb{R}\right\}.$
+Proof 4.29.
+Let $U=\operatorname{co}\{f(t)\mid t\in\mathbb{R}\}$. Define the extended real valued function $I_{U}:\mathbb{R}^{n}\to\mathbb{R}\cup\{-\infty,\infty\}$ as
 $I_{U}(x)=\begin{cases}+\infty,&x\notin U\\ 0,&x\in U\end{cases}.$
@@
 $0\leq I_{U}(F_{\varepsilon}(t))\leq\int_{\operatorname{supp}\varphi_{\varepsilon}}I_{U}(f(t-s))\varphi_{\varepsilon}(s)\mathrm{d}\lambda(s)=0,$
-i.e., $I_{U}(F_{\varepsilon}(t))=0$ so $F_{\varepsilon}(t)\in U$. Since $t$ and $\varepsilon$ are arbitrary, the claim follows. ∎
-Remark 5
-As previously mentioned, it is common that the path is a continuous function defined in a compact subset of $\mathbb{R}$, i.e., $f:[a,b]\to\mathbb{R}^{n}$ with $\infty<a<b<\infty$.
+i.e., $I_{U}(F_{\varepsilon}(t))=0$ so $F_{\varepsilon}(t)\in U$. Since $t$ and $\varepsilon$ are arbitrary, the claim follows.
+Remark 4.30.
+It is common that the path is a continuous function defined in a compact subset of $\mathbb{R}$, i.e., $f:[a,b]\to\mathbb{R}^{n}$ with $\infty<a<b<\infty$.
 In such a case, we can extend the function $f$ to $\mathbb{R}$ as follows to get a new continuous function
@@
 First, we introduce the definition of path length for paths that do not need to be differentiable.
-Definition 5 (Length of $f$)
+Definition 4.31 (Length of $f$).
 Let $f:[a,b]\to\mathbb{R}^{n}$ be a continuous function, and $||\cdot||:\mathbb{R}^{n}\to[0,\infty)$ be any norm in $\mathbb{R}^{n}$.
@@
 […]
-Note that when working with trajectories with compact domain, we must extend them as done in Remark 5.
-Lemma 7
+Note that when working with trajectories with compact domain, we must extend them as done in Remark 4.30.
+Lemma 4.32.
 Let $f:[a,b]\to\mathbb{R}^{n}$ be a continuous function and fix $\varepsilon>0$.
-Let $\bar{f}:[a-\varepsilon,b+\varepsilon]$ be its continuous extension as done in remark 5.
+Let $\bar{f}:[a-\varepsilon,b+\varepsilon]$ be its continuous extension as done in Remark 4.30.
 The following two statements are true:
@@
     If $|t|\leq\varepsilon$ and $g(s)=\bar{f}(s-t)$ for all $s\in[a,b]$, then $L(g)\leq L(\bar{f})=L(f)$.
-Proof:
+Proof 4.33.
 We prove each statement separately.
@@
 Now we are ready for the main result regarding the length of the generated mollified path being shorter or equal than the original.
-Theorem 8
+Theorem 4.34.
 Let $f:[a,b]\to\mathbb{R}^{n}$ be a continuous function and let $\varphi$ be a nonnegative mollifier.
-Fixed $\varepsilon>0$ let $\bar{f}:[a-\varepsilon,b+\varepsilon]\to\mathbb{R}^{n}$ be the continuous extension as in remark 5.
+Fixed $\varepsilon>0$ let $\bar{f}:[a-\varepsilon,b+\varepsilon]\to\mathbb{R}^{n}$ be the continuous extension as in remark 4.30.
 Define $F:[a,b]\to\mathbb{R}^{n}$ as $F=\bar{f}*\varphi_{\varepsilon}$; then $L(F)\leq L(f)$.
-Proof:
+Proof 4.35.
 Take a partition $P$ of $[a,b]$, then
@@
 […]
-thus, by definition of the supremum, $L(F)\leq L(f)$. ∎
-Remark 6
+thus, by definition of the supremum, $L(F)\leq L(f)$.
+Remark 4.36.
 Note that this does not imply that if $f$ is a geodesic between $f(a)$ and $f(b)$, then $F$ is also a geodesic between these points.
 This is because $F(a)\neq f(a)$ or $F(b)\neq f(b)$ may occur.
-While Theorem 6 and Remark 5 guarantee that $F([a,b])=(\bar{f}*\varphi_{\varepsilon})([a,b])\subset\operatorname{co}f([a,b])$, we cannot ensure that $F$ has the same starting and ending points as $f$.
-What Theorem 8 establishes is that by considering the actual starting and ending points of $F$, we can ensure that $L(F)\leq L(f)$.
+While Theorem 4.28 and Remark 4.30 guarantee that $F([a,b])=(\bar{f}*\varphi_{\varepsilon})([a,b])\subset\operatorname{co}f([a,b])$, we cannot ensure that $F$ has the same starting and ending points as $f$.
+What Theorem 4.34 establishes is that by considering the actual starting and ending points of $F$, we can ensure that $L(F)\leq L(f)$.
 An example of this property is shown in Figure 5.
-The original function, which is a linear interpolation of three points in $\mathbb{R}^{2}$ and whose domain is $[a,b]=[0,2]$, is extended to the domain $\varepsilon,2+\varepsilon]$ with $\varepsilon=0.5$ using the extension presented in remark 5.
-As it can be seen from each of its components, $F_{\varepsilon}(a)\neq f(a)$ and $F_{\varepsilon}(b)\neq f(b)$, and clearly $L(F_{\varepsilon})\leq L(f)$ as theorem 8 states.
+The original function, which is a linear interpolation of three points in $\mathbb{R}^{2}$ and whose domain is $[a,b]=[0,2]$, is extended to the domain $\varepsilon,2+\varepsilon]$ with $\varepsilon=0.5$ using the extension presented in remark 4.30.
+As it can be seen from each of its components, $F_{\varepsilon}(a)\neq f(a)$ and $F_{\varepsilon}(b)\neq f(b)$, and clearly $L(F_{\varepsilon})\leq L(f)$ as Theorem 4.34 states.
 Finally note that, while $f_{2}$ can be considered a geodesic between the points $(0,0)$ and $(2,2)$ in $\mathbb{R}^{2}$ using the Euclidean norm, $(f_{2}*\varphi_{\varepsilon})$ is not a geodesic between those two points.
-Figure 5: Visual representation of remark 6.
-The left picture represents the original path $f$ as a blue solid line, and as a red solid line its mollification $F_{\varepsilon}=(f_{1}*\varphi_{\varepsilon},f_{2}*\varphi_{\varepsilon})$ where $\varphi$ is defined in equation (1) and $\varepsilon=0.5$.
+Figure 5: Visual representation of remark 4.36.
+The left picture represents the original path $f$ as a blue solid line, and as a red solid line its mollification $F_{\varepsilon}=(f_{1}*\varphi_{\varepsilon},f_{2}*\varphi_{\varepsilon})$ where $\varphi$ is as in (1) and $\varepsilon=0.5$.
 The middle picture represents the same information but for the first component of the function and its mollification, while the right picture represents the same information but for the second component.
-IV-D The effect of reparametrization and mollification
+4.4 The effect of reparametrization and mollification
 Suppose the desired path is encoded using a continuous function $f:[a,b]\to\mathbb{R}^{n}$ and for a given $\varepsilon>0$ we consider its mollification with parameter $\varepsilon$.
-As we have seen, we first need to extend the function $f$ to $\bar{f}:[a-\varepsilon,b+\varepsilon]$ as is done in Remark 5 or in Theorem 8.
+As we have seen, we first need to extend the function $f$ to $\bar{f}:[a-\varepsilon,b+\varepsilon]$ as is done in Remark 4.30 or in Theorem 4.34.
 It may be of interest to reparametrize the curve so it is normalized, i.e., to find a function $g:[-\varepsilon,1+\varepsilon]\to[a-\varepsilon,b+\varepsilon]$, and consider $\bar{f}\circ g*\varphi_{\varepsilon}:[0,1]\to\mathbb{R}^{n}$, as the desired mollified path.
@@
 […]
-V Curvature guarantees from a sequence of waypoints
+5 Curvature guarantees from a sequence of waypoints
 In this section, we show how to systematically analyze the curvature of the generated path.
@@
 From now on, $\mathbb{R}^{n}$ denotes either $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$.
-V-A The case of three points forming two segments
+5.1 The case of three points forming two segments
 Suppose the desired path can be encoded using a parametric function of the following form.
-Definition 6 (Two line segments function)
+Definition 5.37 (Two line segments function).
 Let $P_{0},P_{1},P_{2}\in\mathbb{R}^{n}$, and let $f:[0,2]\to\mathbb{R}^{n}$ be
@@
 the two-lines segment extended function. Note that $\bar{f}\mid_{[0,2]}=f$. $\square$
+We know by Proposition 4.10 that affine functions are invariant under mollification when Assumption 1 is met and the mollifier is an even function, such as in Example 1.
+From Figure 3 and Proposition 4.13 it is clear that for the curve defined in Definition 5.37 it happens that $f(1)\neq F_{\varepsilon}(1)$ for all $\varepsilon>0$.
+However, it of special interesting to compute the intervals in which the original and the mollified curve coincide, because in those intervals there is no approximation, there is an equality between the original and mollified curve.
+This is easily answered in the next proposition, in which we show that the initial and ending points of the original and mollified curve also coincide.
+Proposition 5.38.
+Let $\bar{f}$ be as in (6).
+Suppose $\varepsilon\in(0,\frac{1}{2})$, define for $r\in\{1,2\}$, $V_{r}:=[r-1+\varepsilon,r-\varepsilon]$ and let $F_{\varepsilon}:=\bar{f}*\varphi_{\varepsilon}$ with $\varphi$ even.
+Then $F_{\varepsilon}|_{V}=\bar{f}|_{V}=f_{V}$, $F_{\varepsilon}(0)=f(0)$, and $F_{\varepsilon}(2)=f(2)$.
+Proof 5.39.
+Let $t\in V_{r}$, and note that for any $s\in[-\varepsilon,\varepsilon]$ it holds that $r-1\leq t-\varepsilon<t-s<t+\varepsilon\leq r$.
+Thus, if $r=1$ the function to be mollified is the line $P_{0}+(P_{1}-P_{0})t$ for any $t\in V_{r}$ and if $r=2$ the function to be mollified is the line $P_{1}+(P_{2}-P_{1})(t-1)$ for any $t\in V_{r}$.
+Thus, by Proposition 4.10 we get the desired result.
+The proof that the initial and ending points are the same it is carried in a similar fashion.
 It is straightforward to note that $f$ is differentiable in $(0,2)\setminus\{1\}$ and $\bar{f}$ in $\mathbb{R}\setminus\{1\}$.
@@
 Note that both $f$ and $\bar{f}$ are continuous functions, hence locally integrable, and from Theorem 1 we have that if $F_{\varepsilon}:=\bar{f}*\varphi_{\varepsilon}$ where $\varphi$ is any mollifier and $\varepsilon>0$ then $F_{\varepsilon}^{\prime}=\bar{f}*\varphi_{\varepsilon}^{\prime}$.
-Fix $t\in\mathbb{R}$, and note that
-$\int_{\mathbb{R}\setminus\{t-1\}}\bar{f}^{\prime}(t-s)\varphi_{\varepsilon}(s)\,\mathrm{d}s$ $=\int_{\mathbb{R}\setminus\{1\}}\bar{f}^{\prime}(s)\varphi_{\varepsilon}(s)\,\mathrm{d}s$
-$=\int_{\mathbb{R}}\bar{f}^{\prime}(s)\varphi_{\varepsilon}(s)\,\mathrm{d}s,$
-because $\lambda(\{1\})=0$, so we could take any function $g$ that is equal to $\bar{f}^{\prime}$ in $\mathbb{R}\setminus\{1\}$ but is defined in $t=1$, and we would reach that $\int_{\mathbb{R}}g(t-s)\varphi_{\varepsilon}(s)=\int_{\mathbb{R}\setminus\{t-1\}}\bar{f}^{\prime}(t-s)\varphi_{\varepsilon}(s)\,\mathrm{d}s$.
-That is, we can evaluate $(\bar{f}^{\prime}*\varphi_{\varepsilon})(1)$ even if $\bar{f}^{\prime}(1)$ is not defined.
-Exploiting the properties of the compact support of the mollifiers
-$\int_{\mathbb{R}}\bar{f}^{\prime}(s)\varphi_{\varepsilon}(t-s)\,\mathrm{d}s$
-$=\bar{f}(s)\varphi_{\varepsilon}(t-s)\mid_{s=t-\varepsilon}^{s=t+\varepsilon}-\int_{\mathbb{R}}\bar{f}(s)(-\varphi_{\varepsilon}^{\prime}(t-s))\,\mathrm{d}s$
-$=\int_{\mathbb{R}}\bar{f}(s)\varphi_{\varepsilon}^{\prime}(t-s)\,\mathrm{d}s=(\bar{f}*\varphi_{\varepsilon}^{\prime})(t)=F_{\varepsilon}^{\prime}(t).$
-This proves that in this case $(\bar{f}^{\prime}*\varphi_{\varepsilon})=(\bar{f}*\varphi_{\varepsilon}^{\prime})=F_{\varepsilon}^{\prime}$, which is a powerful result given that the derivative of $\bar{f}$ is constant almost everywhere.
-Theorem 9
-Let $f:[a,b]\to\mathbb{R}^{n}$ and consider any continuous extension of $f$ to $\mathbb{R}$, i.e., a function $\bar{f}:\mathbb{R}\to\mathbb{R}^{n}$ that is continuous, $\bar{f}\mid_{[a,b]}=f$ and it is differentiable almost everywhere in $\mathbb{R}$.
-If it happens that $f$ is differentiable almost everywhere in $[a,b]$ then it holds that for any $\varepsilon>0$ and mollifier
-[…]
-Proof:
-The proof is immediate following the steps from the previous discussion. For a more general result involving distributions see [16, Sections 6.13 and 6.14]. ∎
-Corollary 1
-Let $f$ be the two lines segment function of equation (5) and $\bar{f}$ its extended version defined in equation (6). Then, for any $\varepsilon>0$ and any mollifier $\varphi$ it holds that
-[…]
-We can now exploit the results of Corollary 1 to obtain a formula for the curvature, as well as an upper bound.
-V-A1 Computing the exact curvature
-We know by the Corollary 1 that $F_{\varepsilon}^{\prime}=(\bar{f}*\varphi^{\prime}_{\varepsilon})=(\bar{f}^{\prime}*\varphi_{\varepsilon})$ everywhere.
+It is not difficult to prove that in this case, $F_{\varepsilon}^{\prime}=\bar{f}*\varphi_{\varepsilon}^{\prime}=\bar{f}^{\prime}*\varphi_{\varepsilon}$ for any $\varepsilon>0$.
+We can now exploit these results to obtain a formula for the curvature, as well as an upper bound.
+5.1.1 Computing the exact curvature
+We know by the discussion above that $F_{\varepsilon}^{\prime}=(\bar{f}*\varphi^{\prime}_{\varepsilon})=(\bar{f}^{\prime}*\varphi_{\varepsilon})$ everywhere.
 Consider a mollifier $\varphi$ and let $\varepsilon>0$.
@@
 Equation (7) is an exact formula for the curvature at each $t\in\mathbb{R}$.
-V-A2 Upper bounding the curvature
+5.1.2 Upper bounding the curvature
 […]
@@
 Figure 6: Left plot represents of a three-point-two-segment function and its mollification.
-Right plot represents the curvature of the function and its upper bound $U_{\kappa}$ which is the right hand side of equation (8).
+Right plot represents the curvature of the function and its upper bound $U_{\kappa}$ which is the right hand side of (8).
 In this case the zone of maximum curvature corresponds to a point really close (or equal) to $t=1$.
@@
 Figure 7: Left plot represents of a three-point-two-segment function and its mollification.
-Right plot represents the curvature of the function and its upper bound $U_{\kappa}$ which is the right hand side of equation (8).
+Right plot represents the curvature of the function and its upper bound $U_{\kappa}$ which is the right hand side of (8).
 In this case the zone of maximum curvature does not correspond to the join of the two segments.
 We have used the mollifier presented in Example 1.
-V-B The general case
+5.2 The general case
 We are going to present how, the natural generalization of the previous computations to a $p>2$ segments curve, gives, in general, a worse result than considering the curvature of each pair of segments locally, and then choosing the most restrictive $\varepsilon_{max}>0$ so that all curvatures are constraint under this $\varepsilon_{max}>0$.
@@
 Suppose we have $p\in\mathbb{N}$ segments with $p\geq 2$.
-Using equation (8) and given a maximum curvature $\kappa_{\max}>0$, we can compute for each pair of consecutive segments its respective [^3] The index $i$ ranges from the first pair to the last pair of consecutive segments.
+Using (8) and given a maximum curvature $\kappa_{\max}>0$, we can compute for each pair of consecutive segments its respective [^1] The index $i$ ranges from the first pair to the last pair of consecutive segments.
 $\varepsilon_{i}>0$ such that, under the three-point-two-segment approximation, their curvatures are upper bounded.
-If $\varepsilon_{i}<\frac{1}{2}$ for all $i$, then equation (8) is exact, because only the two segments used for computing $\varepsilon_{i}$ contribute to the mollification at the junction point.
-In this case, take $\varepsilon=\max_{i}\varepsilon_{i}$, which is valid because the only dependence on $\varepsilon_{i}$ in the right-hand side of equation (8) is through $\frac{1}{\varepsilon_{i}}$; hence, $\varepsilon$ satisfies the bound for each pair of segments.
-If $\varepsilon_{i}>\frac{1}{2}$ for some $i$, then equation (8) becomes an approximation.
+If $\varepsilon_{i}<\frac{1}{2}$ for all $i$, then (8) is exact, because only the two segments used for computing $\varepsilon_{i}$ contribute to the mollification at the junction point.
+In this case, take $\varepsilon=\max_{i}\varepsilon_{i}$, which is valid because the only dependence on $\varepsilon_{i}$ in the right-hand side of (8) is through $\frac{1}{\varepsilon_{i}}$; hence, $\varepsilon$ satisfies the bound for each pair of segments.
+If $\varepsilon_{i}>\frac{1}{2}$ for some $i$, then (8) becomes an approximation.
 In this case, one can either accept an admissible error or use $\varepsilon=\max_{i}\varepsilon_{i}$ as an initial condition for an optimization algorithm that seeks the minimum $\varepsilon>0$ that upper bounds the curvature.
-In either case, equation (8) is a powerful, computationally inexpensive tool that can be used to either compute an exact upper bound for the complete trajectory or reduce computation time in an optimization algorithm.
-VI Numerical validations and real experiments
-To demonstrate the effectiveness of our path generation approach, we present both numerical and experimental results for path following of a mollified path by a unicycle vehicle.
+In either case, (8) is a powerful, computationally inexpensive tool that can be used to either compute an exact upper bound for the complete trajectory or reduce computation time in an optimization algorithm.
+6 Numerical validations and real experiments
+To demonstrate the effectiveness of our path generation approach, we first show a comparison between the path generated by different spline methods and the one by mollification.
+We also present both numerical and experimental results for path following of a mollified path by a unicycle vehicle.
 Specifically, we employ the Singularity-Free Guiding Vector Fields (SF-GVF) path following algorithm [18, 17].
 In brief, SF-GVF takes a parametric path $f\in C^{2}(\mathbb{R},\mathbb{R}^{n})$ as input and constructs a vector field $\chi\in C^{2}(\mathbb{R}^{n},\mathbb{R}^{n})$ whose flow traces the mollified path.
-VI-A Numerical results
-We are going to consider the so called “heart” function as our input path. Define the function
+6.1 Comparison with traditional interpolation methods
+We recall that mollification can generate paths from more general inputs than those for which interpolation is possible, such as the input that considers only $x\in\mathbb{Q}$ in Remark 4.26.
+Nonetheless, to compare traditional interpolation methods with our path generation via mollification, consider a collection of points in $\mathbb{R}^{2}$ that are linearly interpolated, and their $C^{2}$ cubic splines, B-splines and quintic Hermite interpolation splines, i.e, polynomial splines that are twice continuously differentiable and which are computed to specifically pass through the pre-defined collection of points.
+We compare the generated paths between the spline approaches and the proposed mollification in Figure 8.
+Figure 8: Comparison of different approach for path generation.
+The blue line represents the path generated from the linearly interpolated sequence of points shown in black.
+The mollifier used is the one presented in Example 1 with $\varepsilon=0.4$.
+As it can be seen, the mollification approach is the method that resembles the original function the most.
+Moreover, due to Proposition 5.38 we know exactly in which intervals the original function and the mollified one coincide.
+Clearly, this is at the cost of not intersecting the collection of points, except the initial and ending ones.
+Nevertheless, note that numerically speaking, the mollification method is by far the simplest one.
+Indeed, for $C^{2}$ continuity there are several conditions that the coefficients of the spline polynomials must met, and which are extremely sensible to changes in the path.
+That is, a small change in the original path, i.e., a small change in a point of the collection, can generate a big change in the generated path.
+This is not something that happens in mollification, since by Proposition 5.38 and the results of Section 4 we know exactly in which sets the original path and its mollification coincide, where is the mollified curve enclosed, as well as the geometric properties preserved.
+Finally, we want to remark an important fact.
+The polynomials spline here presented are at most $C^{2}$ continuous at the collection of points, and its maximum curvature is not easy to compute between two segments nor the complete spline.
+For this purpose an optimization approach is often carried as mentioned in the introduction.
+In the mollified case, with just a single operation, $C^{\infty}$ continuity is obtained, and its derivatives are easy to compute thanks to Theorem 1.
+We have also provided, for this specific example, an analytical and simple to compute curvature upper bound in (8).
+This implies that mollified path can be computed or changed online in low-cost platforms.
+6.2 Usage with a path following algorithm
+For the path-following part, we are going to consider the so called “heart” function as our input path. Define the function
 $t\in[0,2\pi)\to r(t)=2-2\sin(t)+\sin(t)\frac{\sqrt{|\cos(t)|}}{\sin(t)+1.4}.$
@@
 In both pictures the original path $f$, is shown as a blue dotted line, while as a red dotted line the mollified trajectory $F$ for $\varepsilon_{1}=\varepsilon_{2}=0.4$ in the left picture, and for $\varepsilon_{1}=0.2$ and $\varepsilon_{2}=0.8$ in the right picture.
-The black solid line represents the (now smooth) flow generated by the guiding vector field according to [14] starting from an arbitrary initial position.
+The black solid line represents the (now smooth) flow generated by the guiding vector field according to [17] starting from an arbitrary initial condition (IC).
 A numerical simulation of the vehicle under SF-GVF using $\varepsilon_{1}=\varepsilon_{2}=0.4$ for the “heart” path is shown in the left plot of Figure 9.
@@
 Moreover, as $\varepsilon\to 0$, the trajectory approaches the original path more closely.
-Note that the mollified path lies inside the original path, as predicted by Theorem 6.
+Note that the mollified path lies inside the original path, as predicted by Theorem 4.28.
 In practical terms, this means that the “heart” function can now be used with SF-GVF, extending the applicability of this path following algorithm.
@@
 This is a key advantage of the method: by simply adjusting these parameters, we can ensure that the vehicle follows the curve within its dynamic limits, thereby avoiding issues related to reconverging to the path.
-Moreover, numerical computations show that the length of the original path with respect to the $\ell_{1}$ norm is (in arbitrary units) $25.58$, while the mollified curve has length $23.16$ in the left case and $21.74$ in the right case of Figure 8, as predicted by Theorem 8.
-The same conclusions from Theorem 8 can be verified with respect to any other arbitrary $\ell_{p}$ norm.
+Moreover, numerical computations show that the length of the original path with respect to the $\ell_{1}$ norm is (in arbitrary units) $25.58$, while the mollified curve has length $23.16$ in the left case and $21.74$ in the right case of Figure 9, as predicted by Theorem 4.34.
+The same conclusions from Theorem 4.34 can be verified with respect to any other arbitrary $\ell_{p}$ norm.
 Figure 10: Representation of the three dimensional numerical simulation.
 The original path $f$ is represented as a solid blue line, while the mollified trajectory $F=(F_{k})_{k=1}^{3}=(f_{k}*\varphi_{\varepsilon_{k}})_{k=1}^{3}$ is shown as a dashed red line, with $\varphi$ as in example 1 and $\varepsilon_{k}=1$ for $k\in\{1,2,3\}$.
-The black solid line represents the (now smooth) flow generated by the guiding vector field according to [14] starting from an arbitrary initial position.
+The black solid line represents the (now smooth) flow generated by the guiding vector field according to [17] starting from an arbitrary initial condition (IC).
+6.3 Three dimensional path mollification
 Finally, for completeness, a numerical simulation of a 3D mollified path is presented in Figure 10.
@@
 Indeed, the flow of the guiding vector field converges to the mollified trajectory.
-Moreover, Theorem 8 can also be validated numerically.
+Moreover, Theorem 4.34 can also be validated numerically.
 In this case, the length of the original path in the $\ell_{2}$ norm is (in arbitrary units) $7.38$, while the length of the mollified path is $5.48$.
 An important feature of this approach is its scalability: any trajectory of the form $g:\mathbb{R}\to\mathbb{R}^{n}$ can be mollified component-wise, producing a sufficiently smooth curve that can be further adapted to satisfy a variety of constraints.
-VI-B Experimental results
+6.4 Experimental results
 Before presenting the experimental results, we introduce the software and hardware platforms used in the experiments.
-We also provide the necessary links to the developed software so any interested reader can replicate these experiments.
-VI-B1 Software platform
-We use Paparazzi UAV [18] as a development environment for programming autonomous vehicles.
-Paparazzi is a free and open-source hardware and software project designed to create a flexible autopilot system.
-Although it was initially developed for unmanned aerial vehicles (UAVs), researchers from various universities have extended its applications to other autonomous vehicles.
-Moreover, Paparazzi supports multiple hardware designs, making the interconnection between different modules accessible; the user simply needs to focus on the desired algorithm to implement, e.g., mollification of trajectories.
-Another important component of Paparazzi is its Ground Control Station (GCS), where vehicle positions are visualized and different sensor and internal variables can be monitored via telemetry.
-We have added the necessary code to define, display, and modify the mollification of straight lines on the fly.
-See Figure 10 for a screenshot of the Paparazzi GCS showing the mollification of a linear interpolation of points in $\mathbb{R}^{2}$.
+We also provide the necessary links to the developed software so that any interested reader can replicate these experiments.
+6.4.1 Experimental and Software Platform
+Our experimental platform, shown in Figure 11, is a rover modeled as a unicycle, built around a Matek F765-Wing autopilot with an STM32 microcontroller, integrated IMUs, and support for GNSS, compass, and radio receivers—specifically a Matek M10Q-5883, a Futaba 7008SB receiver, and a Zigbee Xbee telemetry radio.
+The entire system runs on the open-source Paparazzi UAV framework [6], which handles autonomous operation, telemetry, and real-time communication with the Ground Control Station (GCS).
+Through the GCS, the user can issue high-level commands and adjust waypoints on the fly, while the onboard microcontroller recomputes the mollified path in real time whenever a point or parameter is modified, as illustrated in Figure 12.
+Figure 11: Rover vehicle and hardware used during the experiments.
 Figure 12: Capture of an experiment using Paparazzi GCS.
@@
 The orange line represents the trajectory described by the vehicle.
-Using the Paparazzi GCS, the user can issue real-time high-level commands to the Rover’s autopilot via telemetry.
-Examples of such commands include path following instructions for desired waypoints, speed setpoints, and guidance and speed controller constants.
-The rest of the system runs on the vehicle’s onboard computer, enabling fully autonomous operation.
-It is important to note that waypoints can be adjusted to fit the trajectory into the desired area.
-By simply moving the control or endpoints, the user can shape the path as desired.
-Onboard software in a microcontroller recalculates the trajectory by mollifying it using the user-specified parameters whenever a point or parameter is changed.
-This provides considerable flexibility in experimental environments where positional uncertainties may arise.
-VI-B2 Experimental platform
-Figure 11: Rover vehicle and hardware used during the experiments.
-The experimental hardware platform is presented in Figure 11 and can be modeled using the standard unicycle model.
-The rover is equipped with a Matek F765-Wing Autopilot as its central component.
-The autopilot consists of an STM32 microcontroller and integrated Inertial Measurement Units (IMUs) for inertial navigation, along with input/output ports for connecting various sensors including barometers, GNSS, compass, and radio receivers.
-We use the open-source Paparazzi UAV framework for programming and uploading code to the autopilot.
-Paparazzi supports the sensors required for outdoor autonomous operation: GNSS receivers and compass for accurate positioning and yaw attitude, radio receivers as a fail-safe for manual control during algorithmic failures, and telemetry radios enabling ground reception of vehicle states and transmission of commands to the vehicle.
-Our implementation uses the Matek M10Q-5883 GNSS and compass module, a Futaba 7008SB radio receiver, and a Zigbee Xbee telemetry radio.
-VI-B3 Experimental results
-Experimental data extracted from Paparazzi logs is shown in Figure 12.
-We created a continuous but non-differentiable path by linearly interpolating points in $\mathbb{R}^{2}$ as it is done in Section V.
+6.4.2 Experimental data
+We show our logs in Figure 13.
+We created a continuous but non-differentiable path by linearly interpolating points in $\mathbb{R}^{2}$ as it is done in Section 5.
 As noted above, any Paparazzi user can create such trajectories by simply moving points in the ground control station before or during the experiment in real time.
 The experiment uses a curve similar to that in Figure 12.
+Figure 13: Real experiment of path following of a mollified non-differentiable path with a rover vehicle.
+The original non-differentiable path $f$ is in black, the mollified family trajectories $\{F_{\varepsilon_{i}}\}_{i=1}^{6}$ to be followed at each stage are in solid red, its initial point $r_{0}$ is denoted as the blue cross, the position of the vehicle for its corresponding $\varepsilon_{i}>0$ at times $t\in[t_{i-1},t_{i}]$ in a thick solid blue line.
+Finally, the complete trajectory of the vehicle $r(t;r_{0})$ throughout all the stages is the dashed blue line.
 The experimental objective is as follows: Given a desired non-differentiable path $f$, mollify it with parameter $\varepsilon>0$ to obtain $F_{\varepsilon}=(f*\varphi_{\varepsilon})$, where $\varepsilon$ is a function of vehicle speed limited by maximum allowed curvature.
@@
 Note this is merely illustrative; curvature and speed need not be linearly related in practice.
-We adapt the curve to vehicle dynamics as follows: knowing the speed at each segment midpoint and assuming it remains constant until the next segment midpoint, we compute maximum allowed curvature using the linear relationship, then obtain the minimum allowed $\varepsilon>0$ from equation (8).
-This process applied to each segment pair creates a family of curves for different $\varepsilon>0$ values.
-With six segments, we obtain five mollified curves (one per corner) plus one for the default $\varepsilon>0$ value.
-As before, $f:\mathbb{R}\to\mathbb{R}^{2}$ denotes the original path, $\varphi$ is the mollifier from Example 1, $F_{\varepsilon}=(f*\varphi_{\varepsilon})$ is the speed-dependent mollified path, and $r:[0,\infty)\to\mathbb{R}^{2}$ denotes vehicle position as a function of time, with $r(t;r_{0})$ indicating initial position $r_{0}\in\mathbb{R}^{2}$.
-Figure 12 shows the results.
-The family of six mollified paths adapts well to vehicle dynamics, approaching the original path when curvature constraints allow (e.g., segments four and five) while strongly constraining the curve when necessary (e.g., the last two segments).
-The theoretical results hold: mollified curves lie within the convex hull of the original path (Theorem 6), and parameters are sufficiently small that Propositions 2 and 5 apply.
-These experiments validate our approach, demonstrating both theoretical solution of Problem 1 and practical viability on real and affordable hardware.
-Figure 12: Experiments with the real rover vehicle.
-Left column figures represent the original trajectory in black, the mollified family trajectories as red dashed lines, the complete trajectory of the vehicle as a dashed blue line, its initial point as a blue cross, and the position of the vehicle for its corresponding $\varepsilon_{i}>0$ in a thick solid blue line.
-Right column figures represent in a similar fashion, the error between the rover’s position and the path in blue, and the GNSS accuracy in green.
-VII Conclusions
+We are going to consider a curve consisting on seven two dimensional points which are linearly interpolated, as shown in Figure 13.
+The adaptation of the curve to vehicle dynamics operates as follows: at initial time $t_{0}$, we set $\varepsilon_{1}=0.5$ to generate the first mollified curve $F_{\varepsilon_{1}}=f*\varphi_{\varepsilon_{1}}$, where $f$ is the original curve and $\varphi$ is as in Example 1.
+As the vehicle advances and reaches each segment midpoint at times $\{t_{i}\}_{i=1}^{6}$, the speed is measured and used to compute the maximum allowed curvature, from which the minimum permissible $\varepsilon_{i+1}$ is determined via (8).
+This generates a family of six parameters $\{\varepsilon_{i}\}_{i=1}^{6}$ and corresponding mollified curves $\{F_{\varepsilon_{i}}\}_{i=1}^{6}$ that dynamically adapt to vehicle constraints.
+As shown in Figure 13, the resulting paths remain close to the original trajectory where curvature permits, e.g., segments four and five, while strongly constraining the curve when necessary, e.g., the last two segments.
+All theoretical guarantees hold: the mollified curves lie within the convex hull of the original path (Theorem 4.28), and parameters are sufficiently small that Propositions 4.6 and 4.16 apply.
+These experiments validate both the theoretical solution of Problem 1 and the practical viability of our approach on real, affordable hardware.
+7 Conclusions
 In this work, we addressed the problem of efficient path generation to make non-suitable curves, such as linear interpolations from waypoint collections, suitable for path following and trajectory tracking algorithms via mollification.